### v1

# Cost-aware Portfolios in a Large Universe of Assets

Abstract
This paper considers the finite horizon portfolio rebalancing problem in terms of mean-variance optimization, where decisions are made based on current information on asset returns and transaction costs. The study’s novelty is that the transaction costs are integrated within the optimization problem in a high-dimensional portfolio setting where the number of assets is larger than the sample size. We propose portfolio construction and rebalancing models with nonconvex penalty considering two types of transaction cost, the proportional transaction cost and the quadratic transaction cost. We establish the desired theoretical properties under mild regularity conditions. Monte Carlo simulations and empirical studies using S&P 500 and Russell 2000 stocks show the satisfactory performance of the proposed portfolio and highlight the importance of involving the transaction costs when rebalancing a portfolio.
Keywords: High-dimensional Portfolio Optimization, Optimal Rebalancing, SCAD, Transaction Costs.
Acknowledgments: The authors are very thankful to seminar participants at the University of Wisconsin-Madison, Queen Mary University of London, Bristol University, University of California at Davis, and Vanderbilt University.
1 Introduction
Portfolio management attracts great interest in statistics, financial econometrics, and quantitative finance. The mean-variance portfolio allocation model proposed by Markowitz (1952) stands as a cornerstone of modern portfolio theory. The model’s practicality in terms of high dimensionality and trading frictions (e.g., transaction costs) is further studied in this paper.
The classic Markowitz mean-variance formulation considers the following optimization problem in a single period:
$\bm{w}^{*}=\arg\min_{\bm{w}}\bm{w}^{\top}\bm{\Sigma}\bm{w}-\gamma\bm{w}^{\top}\bm{\mu},$ (1.1)
where $\bm{w}$ is the weight vector of the assets in the portfolio, $\gamma$ is the inverse of the risk aversion parameter, $\bm{\mu}=(\mu_{1},\cdots,\mu_{p})^{\top}$ is the mean vector of the excess returns of $p$ assets and $\bm{\Sigma}$ is the covariance matrix of the excess returns. Here we assume that $\sum_{i=1}^{p}w_{i}=1$, so the portfolio is fully-invested on the pool of $p$ risky assets. Samuelson (1975) extends the Markowitz (1952) model, which considers discrete multi-periods portfolio allocation. The above method has an elegant analytic solution that depends only on the excess returns’ mean and covariance matrix. The solution to (1.1) is $\bm{w}^{*}=c_{1}\bm{\Sigma}^{-1}\bm{\mu}-c_{2}\bm{\Sigma}^{-1}\mathbf{1}$, where $c_{1}$ and $c_{2}$ depending only on $\bm{\Sigma}$ and $\bm{\mu}$. However, in modern portfolio management, portfolios commonly exhibit high-dimensional characteristics, introducing new challenges to the traditional framework. Notably, the inverse of the sample covariance matrix is not well defined when the number of assets is larger than the sample size.
Besides the difficulties of covariance matrix estimation in high-dimensional portfolio allocation, transaction costs also make the classical Markowitz mean-variance model sub-optimal. In most empirical studies, transaction costs are retrospectively incorporated by analyzing a portfolio strategy’s performance under given-sized transaction costs. However, in financial practice, the costs of the portfolio are considered during the decision-making process, meaning the costs are a part of the optimization problem. For minor trades that have negligible effects on market prices, transaction cost is generally assumed to be proportional to the trading amount. For larger trades, the literature usually assumes that they have a nonlinear impact on the market price, resulting in quadratic transaction costs (Olivares-Nadal and DeMiguel, 2018). Hautsch and Voigt (2019) illustrated that when incorporating transaction costs into the optimization problem, quadratic transaction costs can be understood as causing the covariance matrix to contract towards a diagonal matrix and a mean shift proportionate to current holdings. Proportional transaction cost, on the other hand, is proportional to the sum of absolute rebalancing, which imposes a stronger penalization on turnover and thus reduces the volume and frequency of rebalancing. Some studies have provided empirical evidence that strategies incorporating transaction fees consistently demonstrate superior performance compared to the minimum variance benchmark methods (Du et al., 2022; Hautsch and Voigt, 2019).
1.1 Main Takeaways
This paper considers the problem faced by a real mutual fund or sophisticated investor. Transaction cost is an integral part of portfolio management, which is passed on to investors who can get returns on the net of such costs. We propose an innovative method for constrained portfolio selection tailored for high-dimensional portfolios, taking into account two different types of transaction costs. Moreover, our approach allows for adaptive weight adjustments in multi-stage scenarios. In other words, we construct the optimal portfolio that considers cost of rebalancing in the optimization procedure.
The main contributions of our work are threefold. Firstly, we propose cost-aware portfolio estimators with nonconvex penalty for portfolio construction and rebalancing, and two types of transaction costs are considered in the proposed estimators. The nonconvex penalization can improve the estimation accuracy in contrast with the $\ell_{1}$ penalization method. We show the theory of the proposed portfolio in the construction and rebalancing stages under both cost functions. Secondly, we adopt the local linear approximation (LLA) algorithm (Zou and Li, 2008) to solve the nonconvex optimization problem and develop the theoretical foundations of the proposed portfolio with LLA algorithm. Thirdly, we introduce data-dependent penalization models for large portfolios in both construction and rebalancing periods that consider transaction costs and demonstrate its superior performance in simulation studies and real data examples.
1.2 Brief Literature Review
Our paper connects to the literature on covariance learning and high-dimensional portfolio optimization. Fan et al. (2015) studied a factor-based risk estimator under many assets and introduced a high-confidence level upper bound to assess the estimation. Lam and Feng (2018) proposed a nonparametrically eigenvalue-regularized integrated covariance matrix estimator. Fan et al. (2019) constructed a robust covariance estimation under the approximate factor model with observed factors. Hafner et al. (2020) developed a Kronecker product model for the covariance matrix estimation and established central limit theorems for such estimator. Ren et al. (2015), Fan and Lv (2016) and Ren et al. (2019) conducted studies on the estimation and statistical inference of sparse precision matrix under the assumption of conditional normal distribution. Nevertheless, practical challenges arise in verifying the normality of high-dimensional data. Wang et al. (2020) introduced a refitted cross-validation estimation method for a high-dimensional precision matrix utilizing its Cholesky decomposition. Other recent works include but are not limited to Engle et al. (2019); Cai et al. (2020); So et al. (2022); Ding et al. (2021). The estimates in these works above converge to the true covariance matrix or precision matrix under certain regularity conditions. However, whether the derived portfolio allocation provides the optimal strategy with cost consideration is still unclear.
Another strand of literature considers various constraints on large portfolio optimization. Jagannathan and Ma (2003) proposed no-short-selling constraints to reduce the portfolio risks. Inspired by Jagannathan and Ma (2003), DeMiguel et al. (2009) imposed the norm constraints to the portfolios. Fan et al. (2012) proposed a gross-exposure constraint method to modify the Markowitz mean-variance optimization problem in a single period to reduce the sensitivity of the resulting allocation to input vectors. Du et al. (2022) showed that under mild conditions, solving the $\ell_{1}$ constrained problem yields the same expected utility as solving the $\ell_{0}$ constrained problem, enabling effective implementation of cardinality constraints through $\ell_{1}$ constraints. Zou (2006); Zhao and Yu (2006); Gai et al. (2013) suggested that $\ell_{1}$ regularization methods demonstrate support recovery property under restrictive regularity conditions while being prone to bias for large coefficients. In contrast, nonconvex penalization methods such as SCAD (Fan and Li, 2001) and MCP (Zhang, 2010) exhibit favorable theoretical properties without relying on restrictive conditions and effectively mitigate the estimation bias associated with Lasso.
On the other hand, the literature on transaction costs and portfolio optimization is way more scarce, especially when large dimensions are considered. One exception is the important work by Ledoit and Wolf (2024), where the authors also consider transaction costs at the optimization stage when constructing mean-variance portfolios. However, their paper focuses only on the empirical part and provides no theoretical guidance. Furthermore, the authors do not consider extra regularization terms to discipline the portfolio in large dimensions. In our opinion, our paper complements the analysis in Ledoit and Wolf (2024).
1.3 Organization of the Paper
The remaining of the article is organized as follows. Section 2 proposes the constrained nonconvex penalization model for high-dimensional portfolios in both construction and rebalacing stages with consideration of transaction cost. We also introduce the LLA to solve the proposed model. In Section 3, we conduct a comprehensive numerical study to examine the performance of the proposed method. Empirical analysis is included in Section 4. We summarize our contribution in Section 5.
Notations: We use the standard notational convention that constants $c_{1}$,$c_{2}$, etc., represent universal positive constants, with values that may change from line to line. We write $f(n)=O(g(n))$ to mean $f(n)\leq cg(n)$ for some constant $c\in(0,\infty)$. For a matrix $\bm{M}$, we write $|||\bm{M}|||_{\infty}$ to denote the $L_{\infty}$ norm, $\|\bm{M}\|_{\max}=$ max$|m_{ij}|$ to denote its entry-wise $L_{\infty}$ norm, and $|||\bm{M}|||_{2}$ to denote the spectral norm. $\lambda_{\min}(\bm{M})$ and $\lambda_{\max}(\bm{M})$ are minimum and maximum eigenvalues of $\bm{M}$. For a vector $\bm{v}=(v_{j})_{j=1}^{p}\in\mathbb{R}^{p}$, we write $\|\bm{v}\|_{k}$, $k=1,2$ to denote its $L_{k}$ norm, $\|\bm{v}\|_{0}$ is the number of nonzero entries of $\bm{v}$, and $L_{\infty}$ norm is $\|\bm{v}\|_{\infty}=$max$|v_{j}|$. For subsets $\mathcal{A}\subset[1,\ldots,p]$ and $\mathcal{B}\subset[1,\ldots,p]$, let $\bm{M}_{\mathcal{A},\mathcal{B}}$ denote the submatrix consisting of the rows indexed by $\mathcal{A}$ and the columns indexed by $\mathcal{B}$, and $\bm{v}_{\mathcal{A}}$ denote the subvector containing elements of $\bm{v}$ indexed by the set $\mathcal{A}$. The cardinality of $\mathcal{A}$ is represented as $|\mathcal{A}|$. The symbol $\odot$ denotes the element-by-element multiplication of two vectors. For a function $h:\mathbb{R}^{p}\rightarrow\mathbb{R}$, we write $\nabla h$ to denote a gradient or subgradient.
2 Methodology
2.1 Cost-aware Portfolio Estimator
We consider an investor who wants to allocate the wealth among $p$ risky assets and assume the asset returns to be stationary. We denote the expected asset returns as
$\bm{\mu}=\left(\mu_{1},\mu_{2},\ldots,\mu_{p}\right)^{\top}.$
Suppose an investor enters the market at day 1 with an initial wealth, allocates the wealth (constructing the first portfolio) at day $n+1$, where $n\in\mathcal{N}^{+}$ is a positive number of trading days between two consecutive rebalances, and reallocates the wealth  [^1] For simplicity but without loss of generality, we assume the investor rebalances the portfolio at same frequency, e.g., monthly, quarterly, annually. at the beginning of each of the following $m-1$ consecutive periods, where $m$ is the total number of portfolio formation decisions. At every decision day $d$, $d=n+1,2n+1,\ldots,mn+1$, we use the historical returns $[\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, that contains observed returns of asset $j$ on its $j$-th column to estimate assets return and risk. Figure 1 illustrates the decision process.
Define the portfolio weight vector at each transaction time point, $t=1,\cdots,m$ as $\bm{w}_{t}$
$\bm{w}_{t}=\left(w_{t,1},w_{t,2},\ldots,w_{t,p}\right)^{\top},$
and the weights are taken at the beginning of period $t$. Considering the transaction cost under the Markowitz mean-variance framework (1.1), the optimal cost-aware portfolio at first construction stage is defined as
$\bm{w}^{*}_{1}:=\mathop{\arg\min}_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\bm{\Sigma}_{1}\bm{w}-\gamma\bm{w}^{\top}\bm{\mu}_{1}+\mathbf{C}_{1}(\bm{w}),$ (2.1)
where $\mathbf{C}_{1}(\bm{w})$ is the transaction cost, depending on the trading activity (changes in portfolio weights). Similarly, the optimal cost-aware portfolio at reallocation stage is defined as
$\bm{\delta}^{*}_{t}:=\mathop{\arg\min}_{\bm{\delta}\in\mathbf{R}^{p},\bm{\delta}^{\top}\mathbf{1}=0}\bm{\delta}^{\top}\bm{\Sigma}_{t}\bm{\delta}-\gamma\bm{\delta}^{\top}\bm{\mu}_{t}+\mathbf{C}_{t}(\bm{\delta}),$ (2.2)
where the weight difference vector $\bm{\delta}_{t}=\bm{w}^{*}_{t}-\bm{w}^{+}_{t-1}$ and $\bm{w}^{+}_{t-1}$ means the pre-rebalancing weight as the value of portfolio changes over the holding period, and $\bm{w}^{+}_{t-1}=(f_{n}\circ f_{n-1}\circ\dots\circ f_{1})(\bm{w}_{t-1})$ for $t\geq 2$, where $f_{i}(\bm{w}):=\frac{\bm{w}\odot(\mathbf{1}+\bm{R}_{(t-2)n+i})}{1+\bm{w}^{\top}\bm{R}_{(t-2)n+i}}$. This method is thus called the cost-aware portfolio estimator (CAPE). This paper proposes a new estimator targeting $\bm{w}^{*}_{t}$ under the high-dimensional model setting.
Figure 1: First portfolio construction and reallocation diagram for multi-periods portfolios selection problems
We assume the true weight vector $\bm{w}_{1}^{*}$ and weight difference vector $\bm{\delta}^{*}_{t}$ are sparse. Denote the support sets of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ as $\mathcal{A}_{1}=\left\{j:w^{*}_{1,j}\neq 0\right\}$ and $\mathcal{A}_{t}=\left\{j:\delta^{*}_{t,j}\neq 0\right\},\ t\geq 2$ , respectively. And the sparsity assumption means the cardinality $s_{t}=|\mathcal{A}_{t}|\ll p$ for $t=1,2,\ldots,m$. We believe this assumption is empirically grounded as many active strategy investors only choose a relatively small number (compared to the vast universe) of assets considering the transaction and management fees. In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to obtain a sparse estimator of $\bm{w}_{1}^{*}$:
$\widehat{\bm{w}}_{1}:$ $=\mathop{\arg\min}_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})+P_{\lambda}(\bm{w}),$ (2.3)
where $\widehat{\bm{\Sigma}}_{1}$ and $\widehat{\bm{\mu}}_{1}$ are based on samples $[\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, $d=n+1$, and $P_{\lambda}(\bm{w})=\sum_{j=1}^{p}P_{\lambda}\left(\bm{w}_{j}\right)$ is the penalty function. The oracle (when the true support set $\mathcal{A}_{1}$ is known) portfolio construction weight estimator is given as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle}}=\left(\widehat{\bm{\bm{w}}}_{\mathcal{A}_{1}}^{\text{oracle}},\mathbf{0}\right)=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}).$
We assume the oracle solution is unique, namely,
$\nabla_{j}\mathcal{L}_{1,t}\left(\widehat{\bm{\bm{w}}}^{\text{oracle}}_{1}\right)+h^{\text{oracle}}=0\quad\forall j\in\mathcal{A}_{1},$ (2.4)
where $h^{\text{oracle}}=\frac{2-\gamma\mathbf{1}_{\mathcal{A}_{1}}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{\mu}_{\mathcal{A}_{1}}}{\mathbf{1}_{\mathcal{A}_{1}}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\mathbf{1}_{\mathcal{A}_{1}}}$ is the Lagrange parameter and $\mathcal{L}_{n,1}(\bm{w})=\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})$.
The above discussion pertains to the construction of oracle estimator $\widehat{\bm{w}}_{1}^{\text{oracle}}$ during the first portfolio construction stage. For the reallocation stage, by replacing $\bm{w}$ with $\bm{\delta}$ and substituting the subscript $1$ with $t$, the explanation extends naturally to the construction of oracle estimator $\widehat{\bm{\delta}}_{t}^{\text{oracle}}$ for the reallocation stage.
A popular choice of $P_{\lambda}(\cdot)$ is the Lasso (Tibshirani, 1996). However, the $\ell_{1}$ penalty is biased, requiring a strong Irrepresentable Condition to be selection consistent (Zhao and Yu, 2006). Nonconvex penalization method may outperform the $\ell_{1}$ method in terms of estimation accuracy and selection consistency. The nonconvex penalty function $P_{\lambda}:\mathbb{R}\rightarrow\mathbb{R}$ satisfies the following conditions:
  • (i)
    $P_{\lambda}(t)$ is increasing and concave in $t\in[0,\infty)$ with $P_{\lambda}(0)=0$;
  • (ii)
    $P_{\lambda}(t)$ is differentiable in $t\in(0,\infty)$ with $P_{\lambda}^{\prime}(0):=P_{\lambda}^{\prime}(0+)\geq a_{1}\lambda$;
  • (iii)
    $P_{\lambda}^{\prime}(t)\geq a_{1}\lambda$ for $t\in\left(0,a_{2}\lambda\right]$;
  • (iv)
    $P_{\lambda}^{\prime}(t)=0$ for $t\in[a\lambda,\infty)$ with the pre-specified constant $a>a_{2}$, where $a_{1}$ and $a_{2}$ are two fixed positive constants.
The mild conditions on $P_{\lambda}(\cdot)$ are typically met by a wide range of regularizers, such as SCAD (Fan and Li, 2001) and MCP (Zhang, 2010). It is easy to verify that $a_{1}=a_{2}=1$ for the SCAD, and its derivative is defined as
$P^{{}^{\prime}}_{\lambda}(|\beta|)=\lambda\left\{I(|\beta|\leq\lambda)+\frac{(a\lambda-|\beta|)_{+}}{(a-1)\lambda}I(|\beta|>\lambda)\right\}$
for some $a>2$, and $P^{{}^{\prime}}_{\lambda}(0+)=\lambda$.
However, the nonconvex nature of the penalty term presents computational challenges due to the potential multiple local optima. To address this issue, Zou and Li (2008) proposed a feasible solution using LLA. This approach transforms the nonconvex penalized model into a sequence of weighted $\ell_{1}$ models, thereby leveraging the computational efficiency of the $\ell_{1}$ method. Zou and Li (2008) suggested using the properly initialized one-step LLA estimator instead of the computationally intensive fully converged estimator. It has been demonstrated that the one-step LLA estimator enjoys the strong oracle property when $\bm{X}^{\top}\bm{X}/n\rightarrow\mathbf{M}$ for some positive definition matrix $\mathbf{M}$. Subsequently, Fan et al. (2014) extended this study to high-dimensional settings and showed that initializing the LLA algorithm with a Lasso solution leads to an oracle solution with high probability. The computation algorithm for first portfolio construction is summarized in Algorithm 1. For the computation algorithm at reallocation stages, it suffices to replace $\bm{w}$ with $\bm{\delta}$ and substitute the subscript $1$ with $t$.
Algorithm 1 Local Linear Approximation (LLA) Algorithm for First Portfolio Construction
1:Initialize $\widehat{\bm{w}}_{1}^{(0)}=\widehat{\bm{w}}_{1}^{\text{initial}}$ and compute the adaptive weight:
$\widehat{\bm{\theta}}^{(0)}=\left(\widehat{\theta}_{1}^{(0)},\ldots,\widehat{\theta}_{p}^{(0)}\right)^{\prime}=\left(P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,1}^{(0)}\right|\right),\ldots,P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,p}^{(0)}\right|\right)\right)^{\prime}$
2:for $k=1,2,...$ do repeat the LLA iteration till convergence
3:     Obtain $\widehat{\bm{w}}_{t}^{(k)}$ by solving the following optimization problem
$\widehat{\bm{w}}_{1}^{(k)}=\min_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\mathcal{L}_{n,1}(\bm{\bm{w}})+\sum_{j}\widehat{\theta}_{j}^{(k-1)}\cdot\left|w_{j}\right|$
4:     Update the adaptive weight vector $\widehat{\bm{\bm{\theta}}}^{(k)}$ with $\widehat{\theta}_{j}^{(k)}=P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,j}^{(k)}\right|\right)$
5:end for
6:Output: The portfolio allocation $\widehat{\bm{w}}_{t}^{\text{LLA}}$
For $t=1,2,\ldots,m$, we assume that the estimators for mean and covariance matrix satisfy the following element-wise max norm deviation condition for some absolute constants $c_{1},c_{2},M_{1},M_{2}$:
$\operatorname{Pr}\left(\|\widehat{\bm{\mu}}_{t}-\bm{\mu}_{t}\|_{\infty}\leq M_{1}\sqrt{\log p/n}\right)\leq 1-O(p^{-c_{1}}),$ (2.5)
$\operatorname{Pr}\left(\|\widehat{\bm{\Sigma}}_{t}-\bm{\Sigma}_{t}\|_{\max}\leq M_{2}\sqrt{\log p/n}\right)\leq 1-O(p^{-c_{2}})$
When the returns $\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}$ are independent and identically distributed sub-Gaussian random vectors with mean $\bm{\mu}$ and covariance $\bm{\Sigma}$, condition 2.5 is satisfied by sample mean and covariance matrix (see Lemma S.1 in the supplementary materials). Moreover, in high-dimensional settings, theoretically refined estimators, such as linear and nonlinear shrinkage estimators, are expected to better satisfy the assumptions outlined above. Notice that we can relax condition 2.5 and allow dependent return data, including strong-mixing ($\alpha$-mixing), $m$-dependent sequence, and physical (or functional) dependence measure for casual time series (Chang et al., 2024). Nonetheless, we do not provide the full theory result of CAPE under dependent return data here for simplicity.
We further assume the estimator of covariance matrix $\widehat{\bm{\Sigma}}_{t}$ satisfies the following restricted strong convexity (RSC) condition:
$\forall\mathbf{\Delta}\in\mathbb{R}^{p},\mathbf{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\mathbf{\Delta}\geq v\|\bm{\Delta}\|_{2}^{2}-\tau\sqrt{\frac{\log p}{n}}\|\bm{\Delta}\|_{1}~{}~{}\text{where}~{}~{}v>0,\tau\geq 0.$ (2.6)
The sample covariance matrix is typically semi-positive definite and $\forall\bm{\Delta}\in\mathbb{R}^{p},\bm{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\Delta}\geq 0$, so the RSC condition hold trivially for $\frac{\|\bm{\Delta}\|_{1}}{\|\bm{\Delta}\|_{2}^{2}}\geq\frac{v}{\tau}\sqrt{\frac{n}{\log p}}$. We only require the condition to hold over a cone of the form $\frac{\|\bm{\Delta}\|_{1}}{\|\bm{\Delta}\|_{2}^{2}}\leq\frac{v}{\tau}\sqrt{\frac{n}{\log p}}$. The RSC condition is widely adopted in high-dimensional estimation and is satisfied with high probability if the return vector is i.i.d. sub-Gaussian random variables (Loh and Wainwright, 2015). To establish the non-asymptotic error bounds, we impose the following assumptions:
  • (A1)
    The estimators $\widehat{\bm{\mu}}_{t}$ and $\widehat{\bm{\Sigma}}_{t}$ the estimators satisfy the element-wise max norm deviation condition (2.5) and the RSC condition (2.6);
  • (A2)
    There exists $C_{1},C_{2}>0$, s.t. $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}\|_{1}<C_{1}$, $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{r}\|_{\infty}<C_{2},\ \forall\bm{r}\in\{-1,0,1\}^{s_{t}}$;
  • (A3)
    The covariance matrix satisfies $0<\lambda_{\min}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})\leq\lambda_{\max}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})<\infty$ and there exists $C_{3}>0$, such that $|||\bm{\Sigma}_{\mathcal{A}^{c}_{t},\mathcal{A}_{t}}|||_{\infty}<C_{3}$;
  • (A4)
    Suppose the theoretical minimizer $\bm{w}_{1}^{*}$ is sparse with cardinality $s_{1}=\|\bm{w}^{*}_{1}\|_{0}$ and the minimal signal strength of $\bm{w}_{1}^{*}$ satisfies that $\|\bm{\bm{w}}_{\mathcal{A}_{1}}^{*}\|_{\min}>(a+1)\lambda$.
Assumptions (A1)–(A3) hold for both the first portfolio construction stage and the reallocation stage, implying that $t=1,2,\ldots,m$. As demonstrated by Liu et al. (2024), the sample mean and covariance matrix satisfy the condition (2.5) in Assumption (A1), if the returns $\bm{R}_{d-n},\ldots,\bm{R}_{d-1}$ are i.i.d. sub-Gaussian random vectors with mean $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$, and $\log p<n$. To improve the estimation accuracy, both linear and nonlinear shrinkage estimator employed in this study are assumed to satisfy condition (2.5). In Assumption (A2), for simplicity, we denote the submatrix of $\bm{\Sigma}_{t}$ consisting of the rows and columns indexed by $\mathcal{A}_{t}$ as $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$, and similarly, the subvector of $\bm{\mu}_{t}$ containing the elements indexed by $\mathcal{A}_{t}$ is represented as $\bm{\mu}_{\mathcal{A}_{t}}$. This notation will be adopted consistently in the subsequent discussion. Assumption (A2) is introduced to ensure that the $\ell_{1}$-norm of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ remain bounded, thereby preventing divergence. Similar assumptions can be found in Cai et al. (2022) and Liu et al. (2024). And the assumption constraining the $\ell_{1}$-norm of the parameter to prevent divergence also appears in Loh and Wainwright (2015). Furthermore, Assumption (A3) is less restrictive than the Irrepresentable Condition required for the Lasso, which restricts the strength of interactions between variables in $\mathcal{A}_{t}$ and $\mathcal{A}^{c}_{t}$. It ensures that the influence of variables outside $\mathcal{A}_{t}$ are not excessively large, preserving sparsity or stability in the covariance structure. Finally, the sparsity assumption in Assumption (A4) is reasonable and has also been employed in Du et al. (2022). Assumption (A4) also imposes the minimal signal strength condition, which is required for selection consistency, and the similar condition can be found in Fan et al. (2014), Loh and Wainwright (2017).
2.2 CAPE with Quadratic Transaction Cost
2.2.1 First Portfolio Construction
We begin with considering the quadratic transaction cost from the initial period, $t=1$,
$\mathbf{C}_{1}(\bm{w})=(\bm{\beta}\odot\bm{w})^{\top}\bm{w}$
where $\bm{\beta}=(\beta_{1},...,\beta_{p})^{\top},\beta_{j}>0$ is the cost parameter, which means the unit transaction cost of the risky asset $j$ is a linear function of the trading size with slope $\beta_{j}$, for $j=1,...,p$. The problem (2.1) then becomes
$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1},\quad\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$
This is a constrained convex optimization problem. By Lagrange multiplier method and the definition of $\bm{w}^{*}$ and $\mathcal{A}_{t}$, we have the explicit solution $\bm{w}_{1}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{1}^{-1}\left(\gamma\bm{\mu}_{1}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{1}^{-1}\bm{\mu}_{1}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{1}^{-1}\mathbf{1}}\mathbf{1}\right)$, and $\bm{w}^{*}_{\mathcal{A}_{1}^{c}}=\mathbf{0}$, by sparsity assumption, where $\tilde{\bm{\Sigma}}_{1}=\bm{\Sigma}_{1}+\mathbb{B}$, $\mathbb{B}=\text{diag}(\beta_{1},...,\beta_{p})$.
As discussed earlier, we use the LLA algorithm for SCAD, and $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ is the initial and Lasso solution. In the following lemma, we show the consistency of the initial Lasso weight estimator.
Lemma 1.
For the first portfolio construction period, suppose that Assumptions (A1)-(A3) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$, where $C_{2}$ is defined in Assumption (A2). Let $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ be the unique optimum of the program
$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}=1}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1}+\lambda_{Lasso}\|\bm{w}\|_{1}$ (2.7)
with $\lambda_{Lasso}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}\sqrt{\frac{\log p}{n}}\right),\quad\quad\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{\frac{s_{1}\log p}{n}}\right).$
Lemma 1 says that the initial Lasso estimator yields an estimate that is consistent in $\ell_{2}$ norm if $\sqrt{s_{1}\log p/n}\to 0$ as $n,p\to\infty$. In other words, if the portfolio has sufficient sparsity that $s_{1}=o(\sqrt{\frac{n}{\log p}})$, the estimation error can be ignored. Our error rate also matches the rate of Lasso estimator for linear regression. For the first portfolio construction, the oracle portfolio construction weight estimator with quadratic transaction cost is given as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{\mathcal{A}_{1}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{w}_{\mathcal{A}_{1}}-\gamma\bm{w}^{\top}_{\mathcal{A}_{1}}\bm{\mu}_{1}+(\bm{\beta}_{{\mathcal{A}_{1}}}\odot\bm{w}_{{\mathcal{A}_{1}}})^{\top}\bm{w}_{\mathcal{A}_{1}}.$ (2.8)
The following theorem states that we can find the oracle estimator after two iterations by implementing the LLA algorithm, and it also provides the convergence rate in $\ell_{\infty}$-norm.
Theorem 1.
Suppose that Assumptions (A1)-(A4) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2), we choose $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$. Let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$. If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$. Furthermore,
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{LLA},\beta}-\bm{w}_{1}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
Theorem 1 states that under mild conditions, our estimator agrees with the oracle estimator $\widehat{\bm{w}}_{1}^{\text{oracle},\beta}$ with high probability when $k$ diverges to infinity. Hence, we only study the property of non-convex penalty in the following sections.
Corollary 1.
Given the same conditions as in Theorem 1, if $s_{1}\sqrt{\log s_{1}/n}=o(1)$, then we have
$|(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{1}-(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{1}|=o_{p}(1),\qquad|(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{2}-(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{2}|=o_{p}(1);$
$|(\widehat{{\bm{w}}}_{1}^{\text{LLA},\beta})^{\top}\widehat{\bm{\Sigma}}_{1}\widehat{\bm{w}}^{\text{LLA},\beta}_{1}-\bm{w}_{1}^{*}\bm{\Sigma}_{1}\bm{w}_{1}^{*}|=o_{p}(1),\qquad|(\widehat{{\bm{w}}}_{1}^{\text{LLA},\beta})^{\top}\widehat{\bm{\Sigma}}_{2}\widehat{\bm{w}}^{\text{LLA},\beta}_{1}-(\bm{w}_{1}^{*})^{\top}\bm{\Sigma}_{2}\bm{w}_{1}^{*}|=o_{p}(1);$
$\left|\frac{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{1}}{\sqrt{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\Sigma}}_{1}\widehat{\bm{\bm{w}}}_{1}}}-\frac{(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{1}}{\sqrt{(\bm{w}_{1}^{*})^{\top}\bm{\Sigma}_{1}\bm{w}_{1}}}\right|=o_{p}(1),\qquad\left|\frac{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\mu}}_{2}}{\sqrt{(\widehat{\bm{w}}^{\text{LLA},\beta}_{1})^{\top}\widehat{\bm{\Sigma}}_{2}\widehat{\bm{\bm{w}}}_{1}}}-\frac{(\bm{w}_{1}^{*})^{\top}\bm{\mu}_{2}}{\sqrt{(\bm{w}_{1}^{*})^{\top}\bm{\Sigma}_{2}\bm{w}_{1}}}\right|=o_{p}(1).$
Corollary 1 shows that if the sparsity level and the dimension satisfy that $s_{1}\sqrt{\log s_{1}/n}=o(1)$, the in-sample and out-of-sample Sharpe ratio estimation error of initial period vanishes when $n$ goes to infinity. Caner et al. (2023) established similar results under a factor model framework. In contrast, we impose a weaker condition on the data-generating process.
2.2.2 Portfolio Reallocation
At reallocation decision day $d$, the transaction costs $C_{t}(\bm{w})$ are a quadratic function of the weight difference vector $\bm{\delta}_{t}$, defined as $\bm{\delta}_{t}:=\bm{w}_{t}-\bm{w}^{+}_{t-1}$. And $\bm{\delta}_{t}$ represents the difference between the new allocation $\bm{w}_{t}$ and the market value-adjusted allocation $\bm{w}^{+}_{t-1}$. The form of transaction costs $C_{t}(\bm{w})$ is thus given by
$\mathbf{C}_{t}(\bm{w})=[\bm{\beta}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})]^{\top}(\bm{w}_{t}-\bm{w}^{+}_{t-1})$
The program (2.1) turns out to be
$\min_{\bm{w}_{t}}\bm{w}^{\top}_{t}\bm{\Sigma}_{t}\bm{w}_{t}-\gamma\bm{w}^{\top}_{t}\bm{\mu}_{t}+[\bm{\beta}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})]^{\top}(\bm{w}_{t}-\bm{w}^{+}_{t-1}),\quad\text{ s.t. }\bm{w}_{t}^{\top}\mathbf{1}=1,$ (2.9)
With $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, the optimization problem has the equivalent form:
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t}+(\bm{\beta}\odot\bm{\delta}_{t})^{\top}\bm{\delta}_{t},\quad\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.10)
Here, we need to modify Assumption (A4) and propose the following revised assumption:
  • (A4’)
    Suppose the theoretical minimizer $\bm{\delta}_{t}^{*}$ is sparse with cardinality $s_{t}=\|\bm{\delta}^{*}_{t}\|_{0}$ and the minimal signal strength of $\bm{\delta}_{t}^{*}$ satisfies that $\|\bm{\bm{\delta}}_{\mathcal{A}_{t}}^{*}\|_{\min}>(a+1)\lambda$;
Similarly, using the Lagrange multiplier method, we have that $\bm{\delta}_{t}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{t}^{-1}(\gamma\tilde{\bm{\mu}}_{t}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{t}^{-1}\tilde{\bm{\mu}}_{t}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{t}^{-1}\mathbf{1}}\mathbf{1})$ and $\bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0}$, where $\tilde{\bm{\mu}}_{t}=\bm{\mu}_{t}-\frac{2}{\gamma}\bm{\Sigma}_{t}\bm{w}^{+}_{t-1}$.
In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to encourage a sparse estimator of $\bm{\delta}_{t}^{*}$
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0,$ (2.11)
where $\widehat{\bm{\Sigma}}_{t}$ and $\widehat{\bm{\mu}}_{t}$ are the estimators of covariance matrix and mean of the asset returns of the t period, $\widehat{\tilde{\bm{\Sigma}}}_{t}=\widehat{\bm{\Sigma}}_{t}+\mathbb{B}$. By incorporating the transaction costs in the mean variance problem and regularizing the eventual rebalancing events, our portfolio achieves better Sharpe ratio and less transaction costs, which is shown in the numerical studies.
Define the oracle estimator (with known true support of $\bm{\delta}$) with quadratic transaction cost under portfolio reallocation as
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=1}\bm{\delta}^{\top}_{\mathcal{A}_{t}}\widehat{\tilde{\bm{\Sigma}}}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}+2(\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t})_{\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}-\gamma\bm{\delta}^{\top}_{\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}.$ (2.12)
Theorem 2.
Suppose that Assumptions (A1)-(A3) and (A4’) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2), let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$ be the unique optimum of the program $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+\lambda_{Lasso,t}\|\bm{\delta}_{t}\|_{1}$. Then we have $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)$ and $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{\frac{s_{t}\log p}{n}}\right)$. If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, we select $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$ and let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso }}.$
  • (i)
    If $\|\bm{w}^{+}_{t-1}\|_{1}<\infty$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log p/n}\right).$
  • (ii)
    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq\infty$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log k/n}\right)$.
Theorem 2 implies that in portfolio reallocation, the LLA algorithm initialized at a Lasso solution yields a local optimum of the nonconvex regularized problem that satisfies oracle properties with low $\ell_{\infty}$ error. Theorem 2 highlights the impact of different estimation methods employed in the previous $t-1$ stages on the performance of our proposed method in stage $t$. Specifically, if regularized estimation methods, such as the proposed CAPE or Lasso, were consistently applied in the previous $t-1$ stages, the resulting $\bm{w}^{+}_{t-1}$ will exhibit a certain level of sparsity. This sparsity enables $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ to achieve a superior convergence rate in the current stage. Conversely, if the estimation methods used in the earlier stages fail to produce sparse estimates, leading to a non-sparse $\bm{w}^{+}_{t-1}$, our method can still ensure that $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\bm{\delta}^{*}_{t}$ with probability approaching 1. However, the convergence rate will be $\sqrt{\log p/n}$, which is attributable to the lack of sparsity in $\bm{w}^{+}_{t-1}$. Corollary 2 provides detailed conclusions regarding the estimation of $\bm{w}^{*}_{t}$ when the proposed CAPE method is consistently applied across the previous $t-1$ stages.
Corollary 2.
Define $\max\{s_{1},\ldots,s_{t}\}\leq k_{t}\leq p$ and $\widehat{\bm{w}}_{t}^{\text{LLA},\beta}=\bm{w}_{t-1}^{+}+\widehat{\bm{\delta}}_{t}^{\text{LLA},\beta}$, and we use CAPE to make estimation in the past $t-1$ stage. Given the same conditions as in Theorem 1 and Theorem 2, with $k_{t}\sqrt{\log k_{t}/n}=o(1)$, then we have
$\|\widehat{\bm{\bm{w}}}_{t}^{\text{LLA},\beta}-\bm{w}_{t}^{*}\|_{\infty}=O_{p}\left(\sqrt{\frac{\log k_{t}}{n}}\right);$
$|(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t}-(\bm{w}_{t}^{*})^{\top}\bm{\mu}_{t}|=o_{p}(1),\qquad|(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t+1}-(\bm{w}_{t}^{*})^{\top}\bm{\mu}_{t+1}|=o_{p}(1);$
$|(\widehat{\bm{w}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}-(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t}\bm{w}_{t}^{*}|=o_{p}(1),\qquad|(\widehat{{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t+1}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}-(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t+1}\bm{w}_{t}^{*}|=o_{p}(1);$
$\left|\frac{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t}}{\sqrt{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}}}-\frac{(\bm{w}_{t}^{*})^{\top}\bm{\mu}_{t}}{\sqrt{(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t}\bm{w}_{t}^{*}}}\right|=o_{p}(1),\ \left|\frac{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\mu}}_{t+1}}{\sqrt{(\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t})^{\top}\widehat{\bm{\Sigma}}_{t+1}\widehat{\bm{\bm{w}}}^{\text{LLA},\beta}_{t}}}-\frac{\bm{w}_{t}^{*}\bm{\mu}_{t+1}}{\sqrt{(\bm{w}_{t}^{*})^{\top}\bm{\Sigma}_{t+1}\bm{w}^{*}_{t}}}\right|=o_{p}(1).$
Corollary 2 shows that if the sparsity level and the dimension satisfy that $k_{t}\sqrt{\log k_{t}/n}=o(1)$, the in-sample and out-of-sample Shape Ratio estimation errors of reallocation periods vanish when $n$ goes to infinity.
2.3 CAPE with Proportional Transaction Cost
2.3.1 First Portfolio Construction
We then consider some real trading scenarios where costs are proportional to the sum of absolute rebalancing ($\ell_{1}$-norm of rebalancing). For the proportional transaction cost, the optimization problem corresponding to the initial period is as follows:
$\min_{\bm{w}_{1}}\bm{w}^{\top}_{1}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$ (2.13)
where $\bm{\alpha}=(\alpha_{1},...,\alpha_{p})^{\top},\alpha_{j}>0$ is the cost parameter, which is similar to $\bm{\beta}$ in Section 2.2. The first-order condition of the constrained optimization (2.13) is
$2\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{\mu}_{1}+\bm{\alpha}\odot\mathbf{g}-h^{*}\mathbf{1}=0,$ (2.14)
where $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{1}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{1}\right\|_{1}/\partial\bm{w}_{1}$, consisting of elements which are 1 or $-1$ in case $w_{j}>0$ or $w_{j}<0$, respectively, or $\in[-1,1]$ in case $w_{j}=0$, and $h^{*}=\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{1}^{-1}\bm{\mu}_{1}+\mathbf{1}^{\top}\bm{\Sigma}_{1}^{-1}(\bm{\alpha}\odot\mathbf{g})}{\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\mathbf{1}}$ denotes the Lagrangian multiplier. By Lagrange multiplier method and the definition of $\bm{w}^{*}_{1}$ and $\mathcal{A}_{1}$, solving for $\bm{w}_{1}^{*}$ yields
$\bm{w}_{1}^{*}=\frac{1}{2}\bm{\Sigma}^{-1}_{t}\left(\gamma\bm{\mu}_{1}-\bm{\alpha}\odot\mathbf{g}+\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}\bm{\mu}_{1}+\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}(\bm{\alpha}\odot\mathbf{g})}{\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{w}^{*}_{\mathcal{A}^{c}_{1}}=\mathbf{0}.$
In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to encourage a sparse estimator of $\bm{w}^{*}_{1}$:
$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+P_{\lambda}(\bm{w}_{1}),\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$ (2.15)
Similar to Lemma 1 in Section 2.2, Lemma 2 shows the consistency of the initial Lasso weight estimator under proportional transaction cost.
Lemma 2.
For the initial period, suppose that Assumptions (A1)-(A3) hold, let $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}$ be the unique optimum of the program
$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}_{1}=1}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+\lambda_{\text{Lasso}}\|\bm{w}_{1}\|_{1},$ (2.16)
with $\lambda_{\text{Lasso}}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$, then with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}(\alpha+\sqrt{\frac{\log p}{n}})\right),\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{s_{1}}(\alpha+\sqrt{\frac{\log p}{n}})\right),$
where $\theta$ is the minimal eigenvalue of $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$ and $\alpha=\|\bm{\alpha}\|_{\infty}$.
Lemma 2 differs from Lemma 1 in Section 2.2 due to the distinct effects of the quadratic and proportional transaction costs. Specifically, the quadratic transaction cost can be interpreted as a perturbation to the diagonal elements of the covariance matrix. In contrast, the proportional transaction cost has a different influence, which is reflected in the convergence rate of $\widehat{\bm{w}}^{\text{Lasso},\alpha}_{1}$. If the convergence rate of $\alpha$ exceeds that of $\lambda_{\text{Lasso}}$, the convergence rate of $\widehat{\bm{w}}^{\text{Lasso},\alpha}_{1}$ will be determined by $\alpha$.
Similar to (2.8), at first portfolio construction stage, we define the oracle portfolio construction weight estimator with proportional transaction cost as
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\alpha}=\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{\mathcal{A}_{1}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{w}_{\mathcal{A}_{1}}-\gamma\bm{w}^{\top}_{\mathcal{A}_{1}}\bm{\mu}_{1}+\|\bm{\alpha}_{\mathcal{A}_{1}}\odot\bm{w}^{\top}_{\mathcal{A}_{1}}\|_{1}.$ (2.17)
Theorem 3.
Under Assumptions (A1)-(A4), we select $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$. Let $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{w}}}_{t}^{\text{Lasso },\alpha}$. If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$ and $\|\bm{\alpha}\|_{\infty}\leq C\sqrt{\frac{\log p}{n}}$ for some large constant $C$, $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{w}}}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$. Furthermore,
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{LLA},\alpha}-\bm{w}_{1}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
To ensure the theoretical properties of subsequent proofs, Theorem 3 reasonably assumes $\|\bm{\alpha}\|_{\infty}=O(\sqrt{\log p/n})$, which matches the convergence rate of $\lambda_{\text{Lasso}}$. This differs from the assumption in Section 2.2, where $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2). This distinction is also evident in our later sections on simulation studies and real data applications, where $\bm{\beta}=0.15\times\mathbf{1}_{p}$ while $\bm{\alpha}=0.001\times\mathbf{1}_{p}$.
Hautsch and Voigt (2019) examined the empirical implications of proportional transaction costs on the Sharpe ratio, demonstrating that smaller values of $\alpha$ yield optimal Sharpe ratios. Our choice of $\alpha$ aligns with this finding, maintaining the same order of magnitude.
Building on the insights from Corollary 1, we derive that under the condition $s_{1}\sqrt{s_{1}\log p}=o(1)$, the estimation errors for in-sample and out-of-sample Sharpe ratios during reallocation periods asymptotically vanish as $n$ approaches infinity. This result highlights the diminishing impact of estimation uncertainty in high-dimensional settings when the sparsity condition is adequately controlled.
2.3.2 Portfolio Reallocation
We model the transaction costs $\mathbf{C}_{t}(\bm{w})$ as a proportional function at dynamic periods given by
$\mathbf{C}_{t}(\bm{w})=\|\bm{\alpha}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})\|_{1}.$
The program (2.1) turns out to be
$\min_{\bm{w}_{t}}\bm{w}^{\top}_{t}\bm{\Sigma}_{t}\bm{w}_{t}-\gamma\bm{w}^{\top}_{t}\bm{\mu}_{t-1}+\|\bm{\alpha}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})\|_{1}\text{ s.t. }\bm{w}_{t}^{\top}\mathbf{1}=1,$
for $t=2,3,...m$, with $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, which has the equivalent form:
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1},\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.18)
Similarly, solving for $\bm{\delta}_{t}^{*}$ yields
$\bm{\delta}_{t}^{*}=\frac{1}{2}\bm{\Sigma}_{t}^{-1}\left(\gamma\tilde{\bm{\mu}}_{t}-\bm{\alpha}\odot\mathbf{g}+\frac{2+\gamma\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\bm{\alpha}\odot\mathbf{g}-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\tilde{\bm{\mu}}_{t}}{\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0},$
where $\tilde{\bm{\mu}}=\bm{\mu}-\frac{2}{\gamma}\bm{\Sigma}_{t}\bm{w}^{+}_{t-1}$ and $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{t}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{t}\right\|_{1}/\partial\bm{w}_{t}$. In practice, the feasible form is
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\widehat{\bm{w}}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.19)
Similar to (2.12), we define the oracle estimator with proportional transaction cost under portfolio reallocation as
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=1}\bm{\delta}^{\top}_{\mathcal{A}_{t}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}+2(\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t})_{\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}-\gamma\bm{\delta}^{\top}_{\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}+\|\bm{\alpha}_{\mathcal{A}_{t}}\odot\bm{\delta}_{\mathcal{A}_{t}}\|_{1}.$ (2.20)
Theorem 4.
Suppose that Assumptions (A1)-(A3) and (A4’) hold and $\|\bm{\alpha}\|_{\infty}=C\sqrt{\frac{\log p}{n}}$ for some large constant $C$. Let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}$ be the unique optimum of the program $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\widehat{\bm{w}}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+\lambda_{Lasso,t}\|\bm{\delta}_{t}\|_{1}$. Then we have $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)$ and $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{s_{t}\frac{\log p}{n}}\right)$. If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, we choose $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$ and let $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}.$
  • (i)
    If $\|\bm{w}^{+}_{t-1}\|_{1}<C_{4}$, $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log p/n}\right).$
  • (ii)
    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq C_{4}$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$, $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log k/n}\right)$.
Remark. Although the conditions and conclusions of Theorem 4 and Theorem 2 appear to be very similar, the differences arise from the distinct assumptions regarding transaction costs. For quadratic transaction costs, we assume only that $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2). In contrast, for proportional transaction costs, we require $\alpha$ to satisfy a specific convergence rate. This distinction stems from the differing impacts of transaction costs on the optimization problem introduced in Section 2.1.
For quadratic transaction costs, it suffices to verify that the modified covariance matrix continues to satisfy our assumptions, as demonstrated in the Lemma LABEL:Property_of_tilde_bSigma provided in the Supplementary Material. However, for proportional transaction costs, Lemma 2 explicitly illustrates the influence of $\alpha$ on both the convergence rate and consistency. This influence persists in Theorems 3 and 4, where the assumptions regarding transaction costs lead to similar conclusions despite the differences in the underlying hypotheses.
Analogous to Corollary 2, it can be established that if the sparsity level and dimensionality satisfy $k_{t}\sqrt{k_{t}\log p}=o(1)$, the estimation errors for both in-sample and out-of-sample Sharpe ratios during reallocation periods converge to zero as $n\to\infty$.
3 Simulation
We follow the methodology outlined by Fan et al. (2008, 2012) to simulate excess returns. Specifically, we assume that the excess return of the $i$th asset is generated from a three-factor model:
$R_{i}=b_{i1}f_{i1}+b_{i2}f_{i2}+b_{i3}f_{i3}+\varepsilon_{i},\quad i=1,\cdots,p.$
First, the factor loadings $\bm{b}_{i}$ are drawn from a trivariate normal distribution $\mathsf{N}\left(\bm{\mu}_{b},\mathbf{cov}_{b}\right)$ with parameters specified in Table 1 below. Once generated, these factor loadings remain fixed throughout the simulations. The returns of the three factors over $n$ periods are also sampled from a trivariate normal distribution $\mathsf{N}\left(\bm{\mu}_{f},\mathbf{cov}_{f}\right)$, with the relevant parameters provided in Table 1. Throughout this simulation study, we assume that $\mathbb{E}(\mathbf{\bm{\epsilon}}\mid\mathbf{f})=\mathbf{0}$ and $\mathsf{Cov}(\bm{\epsilon}\mid\mathbf{f})=\operatorname{diag}\left(\sigma_{1}^{2},\cdots,\sigma_{p}^{2}\right)$. The standard deviations of the idiosyncratic errors $\varepsilon_{i}$ are generated from a gamma distribution with a shape parameter of 3.3586 and a scale parameter of 0.1876. Then, the generated parameters $\left\{\sigma_{i}\right\}$ remain fixed across simulations. Finally, the idiosyncratic errors are generated from a normal distribution, where the standard deviations are set to $\left\{\sigma_{i}\right\}$.
Table 1: Parameters used in the simulation
Parameters for factor loadings
$\bm{\mu}_{b}$ $\mathbf{cov}_{b}$
0.78282 0.029145 0.023873 0.010184
0.51803 0.023873 0.053951 -0.006967
0.41003 0.010184 -0.006967 0.086856
Parameters for factor returns
$\bm{\mu}_{f}$ $\mathbf{cov}_{f}$
0.023558 1.2507 -0.034999 -0.20419
0.012989 -0.034999 0.31564 -0.0022526
0.020714 -0.20419 -0.0022526 0.19303
To examine the performance of the proposed method, we include the following three models for comparison in the simulation study.
  • 1.
    The conventional Markowitz mean-variance model
    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1},\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.1)
    where the $w_{1,i}$ is denoted as the $i$th element of $\bm{w}_{1}$.
  • 2.
    The regularized mean-variance model with $\ell_{1}$ penalty.
    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{t}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\lambda\|\bm{w}_{1}\|_{1},\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.2)
  • 3.
    The mean-variance model with the transaction cost.
    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}_{1}),\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.3)
The models described above are for comparison during the first portfolio construction stage. By replacing $\bm{w}$ with $\bm{\delta}$ and substituting the subscript $1$ with $t$, the corresponding three models for the reallocation stage can be obtained.
We set $p=2000$, $m=5$ and $n=200$, dividing the data into five stages to calculate the out-of-sample annual Return, Cost, Turnover, Leverage and Sharpe ratio. The optimal tuning parameter $\lambda^{opt}$ is selected based on the highest in-sample Sharpe ratio.
The main method is the proposed Cost-Aware Portfolio Estimator (CAPE). For comparison, we utilize three different baseline models: the Mean-Variance (MV) model (3.1), the Penalized Mean-Variance model (PMV) (3.2), and the Cost-Aware Mean-Variance (CMV) model (3.3) to compute portfolio weights at each stage. We employ the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE) used in Ledoit and Wolf (2017) to estimate the covariance matrix $\bm{\Sigma}$. Specifically, we utilize the functions nonlin_shrinkLW() and cov.shrink() from the R packages HDShOP and corpcor to calculate NLSE and LSE, respectively. The results from the NLSE will be presented in the main text, while the results from the LSE are included in the Supplementary Materials.
Table 2: Simulation results with quadratic transaction costs, means and standard errors are calculated based on 1000 replicates, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
Stage Method Return Cost Turnover Leverage SR
S2 MV 8.158(0.273) 3.370(0.004) 16.494(0.010) 7.747(0.005) 1.004(0.034)
PMV 7.911(0.139) 0.740(0.002) 4.363(0.006) 1.681(0.003) 2.020(0.036)
CMV 8.330(0.244) 2.693(0.003) 14.740(0.008) 6.870(0.004) 1.162(0.034)
CAPE 7.881(0.128) 0.616(0.002) 3.855(0.005) 1.428(0.003) 2.196(0.036)
S3 MV 4.138(0.273) 6.482(0.024) 22.684(0.040) 7.754(0.005) 0.436(0.029)
PMV 8.107(0.162) 1.056(0.004) 5.275(0.009) 2.799(0.003) 1.747(0.035)
CMV 9.634(0.246) 4.683(0.015) 19.265(0.029) 6.667(0.004) 1.201(0.031)
CAPE 8.189(0.152) 0.839(0.003) 4.516(0.008) 2.458(0.003) 1.908(0.035)
S4 MV 4.285(0.275) 6.541(0.024) 22.791(0.040) 7.736(0.005) 0.451(0.029)
PMV 8.472(0.173) 1.193(0.004) 5.709(0.011) 3.395(0.003) 1.728(0.036)
CMV 11.748(0.247) 4.564(0.014) 19.036(0.027) 6.527(0.004) 1.476(0.031)
CAPE 8.554(0.162) 0.950(0.004) 4.882(0.009) 3.034(0.003) 1.874(0.036)
S5 MV 4.397(0.271) 6.517(0.024) 22.747(0.041) 7.742(0.005) 0.466(0.029)
PMV 9.036(0.173) 1.244(0.005) 5.902(0.012) 3.733(0.004) 1.804(0.035)
CMV 12.730(0.245) 4.477(0.013) 18.860(0.026) 6.476(0.004) 1.609(0.031)
CAPE 9.137(0.163) 0.994(0.004) 5.057(0.010) 3.369(0.004) 1.954(0.035)
  • Notes: MV, PMV, CMV and CAPE are mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with nonlinear shrinkage estimated covariance matrices), respectively. S1 is the portfolio construction stage, S2-S5 are rebalancing stages, respectively. And the values in columns of Return, Cost, Turnover, and Leverage are in percentages. SR is the Sharpe ratio post transaction costs.
Table 3: Simulation results with proportional transaction costs, means and standard errors are calculated based on 1000 replicates, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
Stage Method Return Cost Turnover Leverage SR
S2 MV 9.879(0.273) 1.649(0.001) 16.494(0.010) 7.747(0.005) 1.284(0.036)
PMV 8.214(0.139) 0.436(0.001) 4.363(0.006) 1.681(0.003) 2.117(0.036)
CMV 9.801(0.268) 1.578(0.001) 15.783(0.010) 7.391(0.005) 1.304(0.036)
CAPE 8.132(0.131) 0.383(0.001) 3.828(0.005) 1.414(0.003) 2.224(0.036)
S3 MV 8.352(0.271) 2.268(0.004) 22.684(0.040) 7.754(0.005) 1.063(0.035)
PMV 8.635(0.162) 0.528(0.001) 5.275(0.009) 2.799(0.003) 1.889(0.036)
CMV 8.448(0.267) 2.171(0.004) 21.713(0.038) 7.610(0.005) 1.092(0.034)
CAPE 8.572(0.155) 0.455(0.001) 4.551(0.008) 2.449(0.003) 1.970(0.036)
S4 MV 8.547(0.275) 2.279(0.004) 22.791(0.040) 7.736(0.005) 1.092(0.035)
PMV 9.094(0.173) 0.571(0.001) 5.709(0.011) 3.395(0.003) 1.890(0.036)
CMV 8.660(0.270) 2.198(0.004) 21.979(0.039) 7.612(0.005) 1.121(0.035)
CAPE 8.994(0.166) 0.494(0.001) 4.945(0.010) 3.037(0.003) 1.947(0.036)
S5 MV 8.639(0.271) 2.275(0.004) 22.747(0.041) 7.742(0.005) 1.105(0.035)
PMV 9.690(0.173) 0.590(0.001) 5.902(0.012) 3.733(0.004) 1.971(0.035)
CMV 8.708(0.270) 2.197(0.004) 21.974(0.040) 7.627(0.005) 1.128(0.035)
CAPE 9.610(0.167) 0.514(0.001) 5.135(0.011) 3.386(0.004) 2.030(0.036)
  • Note: See notes in Table 2.
To emulate real-world investment scenarios, we report the performance of portfolios across four stages (S2-S5). The first stage involves forming the initial portfolio, akin to establishing a new fund at its inception. The second stage corresponds to the first rebalancing, while the third stage entails the second rebalancing, and so forth. Following the conventions established in the literature, we define the parameters for proportional and quadratic transaction costs as $\bm{\beta}=\beta\times\mathbf{1}_{p}$ and $\bm{\alpha}=\alpha\times\mathbf{1}_{p}$, where $\beta=0.15$ and $\alpha=0.001$ respectively. The costs in the following tables are therefore interpreted as the percentage of the normalized initial wealth that needs to cover the costs of transactions. Additionally, we quantify portfolio turnover and leverage at each stage $k=2,3,4,5$. The definitions of these metrics are as follows:
$\text{Turnover}_{k}=\sum\limits_{i=1}^{p}|w_{k,i}-w^{+}_{k-1,i}|,w_{1^{+},i}=0,$
$\text{Leverage}_{k}=\sum\limits_{i=1}^{p}|\min\{w_{k,i},0\}|$
The simulation results are summarized in Tables 2 and 3. The classic Markowitz mean-variance portfolio exhibits the highest turnover, resulting in the largest transaction costs and the lowest out-of-sample Sharpe ratio. In contrast, for both the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE), our proposed method, CAPE, demonstrates the highest Sharpe ratio and the lowest transaction costs. This finding underscores the importance of incorporating costs into the optimization problem.
4 Real Data Application
4.1 S&P 500 Index
We collected daily S&P 500 Index composition stocks data covering the holding period from 02-Jan-2017 to 23-Dec-2020. Our portfolio is rebalanced every 251 trading days, resulting in three decision days and four stages S1-S4 (construction, and then three rebalances). The out-of-sample period extends from 01-Jan-2018 to 23-Dec-2020. On each decision day, the covariance matrix $\bm{\Sigma}$ is estimated using two different methods: the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE), both based on the most recent 251 daily returns. In the main text we show the results using NLSE. The similar results obtained using the LSE are presented in Section S3 of the supplementary materials.
In the study we consider those stocks with complete records over the holding period. Specifically, on the first decision day, a stock has a complete set of 251 days of historical returns, and by the last decision day, it also possesses a full set of returns for the subsequent 251 days. As a result, our stock pool comprises 462 stocks.
In addition to our proposed CAPE method and the three comparison methods (MV, PMV and CMV) shown in the simulation studies, we also include the benchmark equally-weighted portfolio (DeMiguel et al., 2009), where the weights are given as $1/p$. Following the convention of the literature we just label it as ‘1/N’ portfolio in the empirical tables. We assess several performance metrics, including out-of-sample annual return, transaction costs, turnover, leverage, out-of-sample risk, the annualized Sharpe ratio (we report it separately for stages S2-S4) for holding the portfolio at its current rebalanced composition until the next rebalancing, and the overall Sharpe ratio during the whole S2-S4 period. The performances of the various methods are reported in Tables 4 and 5, which present results for proportional and quadratic transaction costs, respectively.
Our proposed CAPE method demonstrates the best overall Sharpe ratio performance during the S2 to S4 phases. In terms of annual returns, the CAPE method has shown the ability to control losses during the period (S2) of overall declines in the S&P 500 and maintaining substantial gains at the lower level of risk in the period of overall profitability (S3-S4). Excluding the equally-weighted (‘1/N’) portfolio, the CAPE method exhibits one of the lowest turnover among actively managed portfolios. More importantly, CAPE strikes a balance between the required return and risk, considering transaction costs. The CAPE is usually the best performer in annual Sharpe ratio, and has the best Sharpe ratio in overall out-of-sample period S2-S4. Furthermore, the CAPE method is more robust than the other comparison methods concerning different transaction cost functions.
Table 4: S&P 500 quadratic transaction costs, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S2 -3.835 0.033 1.000 0.000 -0.244 0.863
S3 29.247 0.002 0.181 0.000 2.336
S4 47.154 0.002 0.153 0.000 1.067
MV S2 -2.797 1.034 4.382 1.691 -0.178 0.883
S3 29.977 2.474 6.754 2.148 2.404
S4 44.761 1.878 5.878 1.755 1.052
PMV S2 -6.943 0.826 3.550 1.275 -0.466 0.704
S3 24.586 0.796 1.859 1.530 2.387
S4 37.516 0.291 0.780 1.460 0.905
CMV S2 -2.189 0.808 3.874 1.437 -0.145 0.755
S3 42.578 3.109 7.611 2.799 2.614
S4 46.451 2.603 6.920 2.385 0.742
CAPE S2 -2.016 0.818 3.444 1.222 -0.133 0.970
S3 30.755 1.183 3.226 1.626 2.822
S4 57.831 0.588 1.842 1.433 1.204
  • Notes: ‘1/N’, MV, PMV, CMV and CAPE are the equally-weighted, mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with nonlinear shrinkage estimated covariance matrices), respectively. S1 is the portfolio construction stage, S2-S4 are rebalancing stages, respectively. And the values in columns of Return and Cost represent percentages. The “overall SR” in the last column is the post-transaction cost Sharpe ratio of the portfolio for holding it from S2-S4 to show the overall performance in our sample period. More details of other performance measure can be found in the notes of Table 2.
Table 5: S&P 500 proportional transaction costs, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S2 -3.903 0.100 1.000 0.000 -0.249 0.862
S3 29.230 0.018 0.181 0.000 2.335
S4 47.141 0.015 0.153 0.000 1.067
MV S2 -2.201 0.438 4.382 1.691 -0.140 0.930
S3 31.775 0.675 6.754 2.148 2.603
S4 46.051 0.588 5.878 1.755 1.083
PMV S2 -6.472 0.355 3.550 1.275 -0.434 0.852
S3 31.187 0.622 6.224 2.140 2.598
S4 58.348 0.535 5.354 1.719 1.103
CMV S2 -7.849 0.429 4.295 1.647 -0.518 0.844
S3 29.955 0.707 7.067 2.299 2.514
S4 63.214 0.634 6.339 1.862 1.148
CAPE S2 -2.270 0.347 3.467 1.234 -0.149 1.000
S3 31.757 0.641 6.410 2.288 2.519
S4 52.784 0.578 5.783 1.802 1.223
  • Note: See notes in Table 4.
4.2 Russel 2000 Index
In this subsection we evaluate the performance of CAPE in the relatively small-cap stocks pool. We collected daily Russell 2000 Index composition stocks data for the holding period from 02-Jan-2017 to 23-Dec-2020. Our portfolio is rebalanced every 251 trading days, resulting in three decision days and four stages. The out-of-sample period extends from 01-Jan-2018 to 23-Dec-2020. We employ the same criterion of complete observations described in Section 4.1, resulting in a pool of 991 stocks. This scenario imposes great challenges for portfolio management in practice, however, it showcases the advantage and applicability of our proposed method. The performance of the various methods is reported in Tables 6 and 7, which correspond to proportional and quadratic transaction costs, respectively.
The performances of MV and CMV methods are notably poor, as demonstrated in Tables 6 and 7. The PMV method exhibits a slightly better overall Sharpe ratio during the S2 to S4 phases compared to the MV and CMV methods. In contrast, our CAPE method significantly outperforms the other approaches, achieving the highest overall Sharpe ratio. The turnover ratio of CAPE is notably the smallest among all actively managed portfolios. Consistent with our findings in the S&P 500 large-cap stocks, the CAPE method demonstrates robustness to variations in transaction cost functions.
Table 6: Russell 2000 quadratic transaction costs, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S2 -8.008 0.015 1.000 0.000 -0.504 0.876
S3 44.196 0.003 0.285 0.000 1.810
S4 140.395 0.015 0.287 0.000 1.247
MV S2 -1.192 0.439 4.176 1.588 -0.080 1.010
S3 133.689 0.706 5.013 1.598 1.113
S4 106.914 1.248 5.681 1.394 1.674
PMV S2 3.114 0.193 1.802 0.401 0.286 1.111
S3 114.843 0.183 1.607 0.713 1.224
S4 102.532 0.332 1.728 0.706 1.583
CMV S2 -0.946 0.413 4.049 1.524 -0.065 1.060
S3 189.224 0.778 5.402 1.716 1.144
S4 188.827 1.518 6.000 1.565 1.562
CAPE S2 8.044 0.154 1.194 0.097 0.865 1.122
S3 116.971 0.089 0.827 0.344 1.278
S4 88.369 0.224 0.833 0.351 1.474
  • Note: See notes in Table 4.
Table 7: Russell 2000 proportional transaction costs, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
Method Stage Return Cost Turnover Leverage SR Overall SR
1/N S2 -8.093 0.100 1.000 0.000 -0.510 0.876
S3 44.171 0.029 0.285 0.000 1.809
S4 140.381 0.029 0.287 0.000 1.247
MV S2 -1.170 0.418 4.176 1.588 -0.079 1.014
S3 133.893 0.501 5.013 1.598 1.115
S4 107.594 0.568 5.681 1.394 1.685
PMV S2 3.127 0.180 1.802 0.401 0.287 1.112
S3 114.865 0.161 1.607 0.713 1.224
S4 102.691 0.173 1.728 0.706 1.585
CMV S2 -3.548 0.439 4.391 1.695 -0.249 1.028
S3 141.701 0.527 5.265 1.735 1.081
S4 126.576 0.745 7.454 2.066 1.841
CAPE S2 8.104 0.119 1.190 0.095 0.872 1.128
S3 116.553 0.082 0.824 0.341 1.280
S4 89.496 0.083 0.833 0.349 1.481
  • Note: See notes in Table 4.
5 Conclusion
In this paper, we propose a novel approach to integrate the transaction cost within the optimization problem when the number of assets is larger than the sample size. By using LLA algorithm, we prove the sign consistency and provide the $\ell_{\infty}$ bound of our proposed CAPE method for both proportional and quadratic transaction costs.
In addition to the sound statistical properties, CAPE demonstrates exceptional performance in practical applications. Through extensive simulations and applications to real data we demonstrate the superior performance and broad applicability of our method across diverse tasks.

### v2

@@
 Abstract
-This paper considers the finite horizon portfolio rebalancing problem in terms of mean-variance optimization, where decisions are made based on current information on asset returns and transaction costs.
-The study’s novelty is that the transaction costs are integrated within the optimization problem in a high-dimensional portfolio setting where the number of assets is larger than the sample size.
-We propose portfolio construction and rebalancing models with nonconvex penalty considering two types of transaction cost, the proportional transaction cost and the quadratic transaction cost.
-We establish the desired theoretical properties under mild regularity conditions.
-Monte Carlo simulations and empirical studies using S&P 500 and Russell 2000 stocks show the satisfactory performance of the proposed portfolio and highlight the importance of involving the transaction costs when rebalancing a portfolio.
-Keywords: High-dimensional Portfolio Optimization, Optimal Rebalancing, SCAD, Transaction Costs.
+This paper proposes a finite-horizon (multi-period) mean-variance portfolio estimator, in which rebalancing decisions are based on current information about asset returns and transaction costs.
+The novelty of this study stems from integrating transaction costs into the decision process within a high-dimensional portfolio setting, where the number of assets exceeds the sample size.
+We define the optimal cost-aware portfolio and propose novel models for its construction and rebalancing.
+Our approach incorporates a nonconvex penalty and explicitly accounts for both proportional and quadratic transaction costs.
+We establish that the estimators derived from the proposed construction and rebalancing models, as well as the corresponding in-sample and out-of-sample Sharpe ratio estimators, consistently converge to those of the optimal cost-aware portfolio.
+Monte Carlo simulations and empirical studies using S&P 500 and Russell 2000 stocks show the satisfactory performance of the proposed portfolio and highlight the importance of incorporating transaction costs during rebalancing.
+Keywords: High-dimensional Portfolio Optimization, Mean-variance Model, Optimal Rebalancing, Transaction Costs.
 Acknowledgments: The authors are very thankful to seminar participants at the University of Wisconsin-Madison, Queen Mary University of London, Bristol University, University of California at Davis, and Vanderbilt University.
@@
 Here we assume that $\sum_{i=1}^{p}w_{i}=1$, so the portfolio is fully-invested on the pool of $p$ risky assets.
-Samuelson (1975) extends the Markowitz (1952) model, which considers discrete multi-periods portfolio allocation.
-The above method has an elegant analytic solution that depends only on the excess returns’ mean and covariance matrix.
-The solution to (1.1) is $\bm{w}^{*}=c_{1}\bm{\Sigma}^{-1}\bm{\mu}-c_{2}\bm{\Sigma}^{-1}\mathbf{1}$, where $c_{1}$ and $c_{2}$ depending only on $\bm{\Sigma}$ and $\bm{\mu}$.
+Samuelson (1975) extends the Markowitz (1952)model to consider discrete multi-period portfolio allocation.
+This approach has an elegant analytical solution that depends only on the mean and covariance matrix of the excess returns.
+The solution to (1.1) is $\bm{w}^{*}=c_{1}\bm{\Sigma}^{-1}\bm{\mu}-c_{2}\bm{\Sigma}^{-1}\mathbf{1}$, where $c_{1}$ and $c_{2}$ depend only on $\bm{\Sigma}$ and $\bm{\mu}$.
 However, in modern portfolio management, portfolios commonly exhibit high-dimensional characteristics, introducing new challenges to the traditional framework.
@@
 In most empirical studies, transaction costs are retrospectively incorporated by analyzing a portfolio strategy’s performance under given-sized transaction costs.
-However, in financial practice, the costs of the portfolio are considered during the decision-making process, meaning the costs are a part of the optimization problem.
-For minor trades that have negligible effects on market prices, transaction cost is generally assumed to be proportional to the trading amount.
+However, in financial practice, the costs of the portfolio are considered during the decision-making process, meaning the costs are an integral part of the rebalancing.
+Transaction costs are generally assumed to be proportional to the trading amount for minor trades with negligible market impact.
 For larger trades, the literature usually assumes that they have a nonlinear impact on the market price, resulting in quadratic transaction costs (Olivares-Nadal and DeMiguel, 2018).
 Hautsch and Voigt (2019) illustrated that when incorporating transaction costs into the optimization problem, quadratic transaction costs can be understood as causing the covariance matrix to contract towards a diagonal matrix and a mean shift proportionate to current holdings.
-Proportional transaction cost, on the other hand, is proportional to the sum of absolute rebalancing, which imposes a stronger penalization on turnover and thus reduces the volume and frequency of rebalancing.
+Proportional transaction costs scale with the sum of the absolute values of rebalancing amounts, penalizing turnover more heavily and thereby reducing rebalancing volume and frequency.
 Some studies have provided empirical evidence that strategies incorporating transaction fees consistently demonstrate superior performance compared to the minimum variance benchmark methods (Du et al., 2022; Hautsch and Voigt, 2019).
@@
 This paper considers the problem faced by a real mutual fund or sophisticated investor.
-Transaction cost is an integral part of portfolio management, which is passed on to investors who can get returns on the net of such costs.
+Transaction costs are an integral part of portfolio management and are passed on to investors, who receive returns net of such costs.
 We propose an innovative method for constrained portfolio selection tailored for high-dimensional portfolios, taking into account two different types of transaction costs.
 Moreover, our approach allows for adaptive weight adjustments in multi-stage scenarios.
-In other words, we construct the optimal portfolio that considers cost of rebalancing in the optimization procedure.
+Specifically, at each decision point, the portfolio weights are rebalanced in a data-driven manner based on the current holdings and the return information from the previous stage.
+In other words, we construct the optimal portfolio that considers the cost of rebalancing, which provides guidance to cost-aware investors.
 The main contributions of our work are threefold.
-Firstly, we propose cost-aware portfolio estimators with nonconvex penalty for portfolio construction and rebalancing, and two types of transaction costs are considered in the proposed estimators.
-The nonconvex penalization can improve the estimation accuracy in contrast with the $\ell_{1}$ penalization method.
-We show the theory of the proposed portfolio in the construction and rebalancing stages under both cost functions.
+Firstly, we propose cost-aware optimal portfolio estimators for new construction and subsequent rebalancing in a high-dimensional framework.
+To our knowledge, such a portfolio estimator in high dimensions appears to be novel.
+Also, the nonconvex penalization employed in our approach can improve the estimation accuracy compared to the $\ell_{1}$ penalization method.
+We establish theoretical properties for the proposed portfolio in both the construction and rebalancing stages under each cost function.
+We show that, under both cost functions, the estimators obtained in the construction and rebalancing stages—as well as the corresponding in-sample and out-of-sample Sharpe ratio estimators—consistently converge to those of the optimal cost-aware portfolio.
 Secondly, we adopt the local linear approximation (LLA) algorithm (Zou and Li, 2008) to solve the nonconvex optimization problem and develop the theoretical foundations of the proposed portfolio with LLA algorithm.
-Thirdly, we introduce data-dependent penalization models for large portfolios in both construction and rebalancing periods that consider transaction costs and demonstrate its superior performance in simulation studies and real data examples.
+Thirdly, we introduce data-dependent penalization for investors facing a large universe of assets in each period while considering transaction costs.
+We demonstrate the superior performance of the proposed portfolio through extensive simulation studies and empirical evaluations.
+In particular, we compare it against a wide range of existing portfolio construction methods.
+For the real data analysis, we use stock return data from the S&P 500 and Russell 2000 indices over the period 2017–2020.
 1.2 Brief Literature Review
@@
 (2021).
-The estimates in these works above converge to the true covariance matrix or precision matrix under certain regularity conditions.
+To enhance out-of-sample performance, Luo et al.
+(2024) proposed an iterative optimization framework based on almost second-degree stochastic dominance (ASSD), which improves convergence efficiency and empirically yields higher out-of-sample returns compared to traditional dominance and mean-variance approaches.
+The estimators in the works above converge to the true covariance matrix or precision matrix under certain regularity conditions.
 However, whether the derived portfolio allocation provides the optimal strategy with cost consideration is still unclear.
@@
 (2009) imposed the norm constraints to the portfolios.
+In particular, they developed a unified framework for portfolio optimization by constraining the norm of the portfolio weights, which effectively mitigates estimation error and encompasses several existing strategies such as shrinkage-based estimators and the 1/N portfolio.
 Fan et al.
@@
 Zou (2006); Zhao and Yu (2006); Gai et al.
-(2013) suggested that $\ell_{1}$ regularization methods demonstrate support recovery property under restrictive regularity conditions while being prone to bias for large coefficients.
+(2013) suggested that $\ell_{1}$ regularization methods demonstrate the support recovery property under restrictive regularity conditions while being prone to bias for large coefficients.
 In contrast, nonconvex penalization methods such as SCAD (Fan and Li, 2001) and MCP (Zhang, 2010) exhibit favorable theoretical properties without relying on restrictive conditions and effectively mitigate the estimation bias associated with Lasso.
-On the other hand, the literature on transaction costs and portfolio optimization is way more scarce, especially when large dimensions are considered.
+In addition, Ban et al.
+(2018) proposed performance-based regularization and cross-validation techniques for portfolio optimization, casting the problem into a robust optimization framework and showing that their method outperforms traditional and regularized approaches across several Fama-French datasets.
+On the other hand, the literature on transaction costs and portfolio optimization in large dimensions is significantly less abundant.
+Most of the papers in Mathematical Finance and Operations Research focus on low-dimensional settings.
+See, for example, Lobo et al.
+(2007) and Olivares-Nadal and DeMiguel (2018).
 One exception is the important work by Ledoit and Wolf (2025), where the authors also consider transaction costs at the optimization stage when constructing mean-variance portfolios.
@@
 In our opinion, our paper complements the analysis in Ledoit and Wolf (2025).
+From an algorithmic perspective, Best and Hlouskova (2005) propose an efficient method for portfolio optimization with transaction costs, which reduces a high-dimensional constrained problem into a sequence of lower-dimensional subproblems by implicitly incorporating transaction cost effects.
+Similarly, Perold (1984) provides a computationally efficient algorithm for large-scale mean-variance portfolio optimization with transaction costs and factor-based models, enabling parametric solutions under practical resource constraints.
+In the dynamic setting, Brown and Smith (2011) study multi-period portfolio optimization with transaction costs and propose heuristic strategies that perform close to optimal, supported by dual-based upper bounds derived from information relaxation.
 1.3 Organization of the Paper
-The remaining of the article is organized as follows.
+The remainder of the article is organized as follows.
 Section 2 proposes the constrained nonconvex penalization model for high-dimensional portfolios in both construction and rebalacing stages with consideration of transaction cost.
@@
 We summarize our contribution in Section 5.
-Notations: We use the standard notational convention that constants $c_{1}$,$c_{2}$, etc., represent universal positive constants, with values that may change from line to line.
-We write $f(n)=O(g(n))$ to mean $f(n)\leq cg(n)$ for some constant $c\in(0,\infty)$.
-For a matrix $\bm{M}$, we write $|||\bm{M}|||_{\infty}$ to denote the $L_{\infty}$ norm, $\|\bm{M}\|_{\max}=$ max$|m_{ij}|$ to denote its entry-wise $L_{\infty}$ norm, and $|||\bm{M}|||_{2}$ to denote the spectral norm.
-$\lambda_{\min}(\bm{M})$ and $\lambda_{\max}(\bm{M})$ are minimum and maximum eigenvalues of $\bm{M}$.
-For a vector $\bm{v}=(v_{j})_{j=1}^{p}\in\mathbb{R}^{p}$, we write $\|\bm{v}\|_{k}$, $k=1,2$ to denote its $L_{k}$ norm, $\|\bm{v}\|_{0}$ is the number of nonzero entries of $\bm{v}$, and $L_{\infty}$ norm is $\|\bm{v}\|_{\infty}=$max$|v_{j}|$.
-For subsets $\mathcal{A}\subset[1,\ldots,p]$ and $\mathcal{B}\subset[1,\ldots,p]$, let $\bm{M}_{\mathcal{A},\mathcal{B}}$ denote the submatrix consisting of the rows indexed by $\mathcal{A}$ and the columns indexed by $\mathcal{B}$, and $\bm{v}_{\mathcal{A}}$ denote the subvector containing elements of $\bm{v}$ indexed by the set $\mathcal{A}$.
-The cardinality of $\mathcal{A}$ is represented as $|\mathcal{A}|$.
-The symbol $\odot$ denotes the element-by-element multiplication of two vectors.
-For a function $h:\mathbb{R}^{p}\rightarrow\mathbb{R}$, we write $\nabla h$ to denote a gradient or subgradient.
+Notations: We use the standard convention that constants $c_{1}$, $c_{2}$, etc., denote universal positive constants whose values may vary from line to line.
+We write $f(n)=O(g(n))$ if $f(n)\leq cg(n)$ for some constant $c>0$.
+For a matrix $\bm{M}$, we let $|\bm{M}|||_{\infty}$ denote the $L_{\infty}$ operator norm, $\|\bm{M}\|_{\max}=\max|m_{ij}|$ denote the entrywise $L_{\infty}$ norm, and $|\bm{M}|||_{2}$ denote the spectral norm.
+The minimal and maximal eigenvalues of $\bm{M}$ are denoted by $\lambda_{\min}(\bm{M})$ and $\lambda_{\max}(\bm{M})$, respectively.
+For a vector $\bm{v}=(v_{j})_{j=1}^{p}\in\mathbb{R}^{p}$, we use $\|\bm{v}\|_{k}$ ($k=1,2$) for its $L_{k}$ norm, $\|\bm{v}\|_{0}$ for the number of nonzero entries, and $\|\bm{v}\|_{\infty}=\max|v_{j}|$ for the $L_{\infty}$ norm.
+Given index sets $\mathcal{A},\mathcal{B}\subset\{1,\ldots,p\}$, we write $\bm{M}_{\mathcal{A},\mathcal{B}}$ for the submatrix of $\bm{M}$ with rows indexed by $\mathcal{A}$ and columns by $\mathcal{B}$, and $\bm{v}_{\mathcal{A}}$ for the subvector of $\bm{v}$ with entries in $\mathcal{A}$.
+The cardinality of $\mathcal{A}$ is denoted $|\mathcal{A}|$.
+The symbol $\odot$ represents element-wise multiplication.
+For a function $\mathbb{R}^{p}\rightarrow\mathbb{R}, h\nabla h$ denotes its gradient or subgradient.
+To simplify notation, when referring to time-specific quantities such as $\bm{\mu}_{t}$ and $\bm{\Sigma}_{t}$, we write $\bm{\mu}_{\mathcal{A}_{t}}$ and $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$ in place of the more precise $\bm{\mu}_{t,\mathcal{A}_{t}}$ and $\bm{\Sigma}_{t,\mathcal{A}_{t},\mathcal{A}_{t}}$ whenever the context is clear.
 2 Methodology
+2.1 Preliminary
+To contextualize our proposed approach, we review three foundational models in the mean-variance portfolio optimization literature.
+These models serve as natural precursors to our method and are also employed as benchmark comparisons in our simulation studies.
+  • 1.
+    The classical Markowitz mean-variance model.
+Originally introduced by Markowitz (1952), this model aims to balance risk and return by minimizing portfolio variance for a given level of expected return.
+It is formulated as:
+    $\min_{\bm{w}}\ \bm{w}^{\top}\widehat{\bm{\Sigma}}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}},\quad\text{subject to }\sum_{i=1}^{p}w_{i}=1,$ (2.1)
+    where $w_{i}$ denotes the $i$th element of the portfolio weight vector $\bm{w}$, $\widehat{\bm{\Sigma}}$ is the estimated covariance matrix, and $\widehat{\bm{\mu}}$ is the vector of estimated expected returns.
+Despite its theoretical elegance, the Markowitz model often suffers from poor empirical performance in high-dimensional settings due to estimation errors in the covariance matrix.
+  • 2.
+    The penalized mean-variance model with gross-exposure constraints.
+Fan et al.
+(2012) propose a regularized version of the classical mean-variance framework by incorporating an $\ell_{1}$ penalty on the portfolio weights to control gross exposure and improve stability in high-dimensional settings.
+The optimization problem is:
+    $\min_{\bm{w}}\ \bm{w}^{\top}\widehat{\bm{\Sigma}}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}+\lambda\|\bm{w}\|_{1},\quad\text{subject to }\sum_{i=1}^{p}w_{i}=1,$ (2.2)
+    where $\lambda>0$ is a tuning parameter that regulates the portfolio’s gross exposure.
+This formulation mitigates the accumulation of estimation errors in large portfolios and can outperform no-short-sale constraints by allowing limited short positions.
+  • 3.
+    The mean-variance model with transaction costs.
+Hautsch and Voigt (2019) extend the mean-variance paradigm to account for transaction costs explicitly.
+By penalizing turnover directly, their model promotes more stable allocations while implicitly regularizing the covariance matrix.
+The corresponding optimization problem is:
+    $\min_{\bm{w}}\ \bm{w}^{\top}\widehat{\bm{\Sigma}}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}+\mathbf{C}(\bm{w}),\quad\text{subject to }\sum_{i=1}^{p}w_{i}=1,$ (2.3)
+    where $\mathbf{C}(\bm{w})$ denotes the transaction cost function, often proportional or quadratic in the change of portfolio weights.
+This model demonstrates that incorporating transaction costs ex ante leads to empirically superior and more stable portfolios, especially under model uncertainty.
+Each of these models contributes essential insights: the classical mean-variance framework provides the theoretical foundation, the penalized model addresses high-dimensional estimation risk, and the cost-aware model introduces practical considerations arising in real-world trading environments.
+Our proposed approach, cost-aware portfolio estimator (CAPE), builds upon these developments by integrating both regularization and dynamic transaction cost modeling in a unified framework, tailored for large-scale portfolio optimization.
 2.2 Cost-aware Portfolio Estimator
-We consider an investor who wants to allocate the wealth among $p$ risky assets and assume the asset returns to be stationary. We denote the expected asset returns as
-$\bm{\mu}=\left(\mu_{1},\mu_{2},\ldots,\mu_{p}\right)^{\top}.$
-Suppose an investor enters the market at day 1 with an initial wealth, allocates the wealth (constructing the first portfolio) at day $n+1$, where $n\in\mathcal{N}^{+}$ is a positive number of trading days between two consecutive rebalances, and reallocates the wealth  [^1] For simplicity but without loss of generality, we assume the investor rebalances the portfolio at same frequency, e.g., monthly, quarterly, annually.
-at the beginning of each of the following $m-1$ consecutive periods, where $m$ is the total number of portfolio formation decisions.
-At every decision day $d$, $d=n+1,2n+1,\ldots,mn+1$, we use the historical returns $[\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, that contains observed returns of asset $j$ on its $j$-th column to estimate assets return and risk.
+We consider an investor who aims to allocate wealth across $p$ risky assets over $m$ periods, assuming that asset returns are stationary within each period $t=1,\ldots,m$.
+We denote the expected asset returns and the corresponding population covariance matrix in the $t$-th period by $\bm{\mu}_{t}$ and $\bm{\Sigma}_{t}$, respectively.
+Suppose an investor enters the market at day 1 with an initial wealth, allocates the wealth (constructing the first portfolio) at day $n+1$, where $n\in\mathcal{N}^{+}$ is a fixed number of trading days between two consecutive rebalances, and reallocates the wealth  [^1] For simplicity but without loss of generality, we assume the investor rebalances the portfolio at same frequency, e.g., annually as in Section 4.
+at the beginning of each of the following m1m-1 consecutive periods, where $m$ is the total number of portfolio formation decisions.
+At every decision day $d$, $d=n+1,2n+1,\ldots,mn+1$, we use the historical returns $\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, that contains observed returns of asset $j$ on its $j$-th column to estimate assets return and risk.
 Figure 1 illustrates the decision process.
@@
 and the weights are taken at the beginning of period $t$.
-Considering the transaction cost under the Markowitz mean-variance framework (1.1), the optimal cost-aware portfolio at first construction stage is defined as
-$\bm{w}^{*}_{1}:=\mathop{\arg\min}_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\bm{\Sigma}_{1}\bm{w}-\gamma\bm{w}^{\top}\bm{\mu}_{1}+\mathbf{C}_{1}(\bm{w}),$ (2.1)
-where $\mathbf{C}_{1}(\bm{w})$ is the transaction cost, depending on the trading activity (changes in portfolio weights).
-Similarly, the optimal cost-aware portfolio at reallocation stage is defined as
-$\bm{\delta}^{*}_{t}:=\mathop{\arg\min}_{\bm{\delta}\in\mathbf{R}^{p},\bm{\delta}^{\top}\mathbf{1}=0}\bm{\delta}^{\top}\bm{\Sigma}_{t}\bm{\delta}-\gamma\bm{\delta}^{\top}\bm{\mu}_{t}+\mathbf{C}_{t}(\bm{\delta}),$ (2.2)
-where the weight difference vector $\bm{\delta}_{t}=\bm{w}^{*}_{t}-\bm{w}^{+}_{t-1}$ and $\bm{w}^{+}_{t-1}$ means the pre-rebalancing weight as the value of portfolio changes over the holding period, and $\bm{w}^{+}_{t-1}=(f_{n}\circ f_{n-1}\circ\dots\circ f_{1})(\bm{w}_{t-1})$ for $t\geq 2$, where $f_{i}(\bm{w}):=\frac{\bm{w}\odot(\mathbf{1}+\bm{R}_{(t-2)n+i})}{1+\bm{w}^{\top}\bm{R}_{(t-2)n+i}}$.
-This method is thus called the cost-aware portfolio estimator (CAPE).
-This paper proposes a new estimator targeting $\bm{w}^{*}_{t}$ under the high-dimensional model setting.
+In real-world applications, investors tend to allocate wealth across a limited number of assets due to transaction costs, regulatory constraints, or interpretability concerns.
+To incorporate this preference, we consider the transaction cost under the Markowitz mean-variance framework (1.1).
+And the optimal cost-aware portfolio at first portfolio construction stage ($t=1$) is defined as
+$\bm{w}^{*}_{1}:=\mathop{\arg\min}_{\bm{w}\in\mathbb{R}^{p}}\bm{w}^{\top}\bm{\Sigma}_{1}\bm{w}-\gamma\bm{w}^{\top}\bm{\mu}_{1}+\mathbf{C}_{1}(\bm{w})\quad\text{s.t.}\ \bm{w}^{\top}\textbf{1}=1,\ \|\bm{w}\|_{0}\leq s_{0},$ (2.4)
+where $\mathbf{C}_{1}(\bm{w})$ is the transaction cost at first portfolio construction stage stage ($t=1$), depending on the trading activity (changes in portfolio weights) and $s_{0}<\infty$ denotes the maximum number of assets allowed in the portfolio.
+Similarly, the optimal cost-aware portfolio at reallocation stage ($t\geq 2$) is defined as
+$\bm{\delta}^{*}_{t}:=\mathop{\arg\min}_{\bm{\delta}\in\mathbb{R}^{p}}\bm{\delta}^{\top}\bm{\Sigma}_{t}\bm{\delta}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}\bm{\mu}_{t}+\mathbf{C}_{t}(\bm{\delta})\quad\text{s.t.}\ \bm{\delta}^{\top}\textbf{1}=0,\ \|\bm{\delta}\|_{0}\leq s_{0},$ (2.5)
+where the weight difference vector $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, $\mathbf{C}_{t}(\bm{w})$ is the transaction cost at reallocation stage ($t\geq 2$), and $\bm{w}^{+}_{t-1}$ means the pre-rebalancing weight as the value of portfolio changes over the holding period, and $\bm{w}^{+}_{t-1}=(f_{n}\circ f_{n-1}\circ\dots\circ f_{1})(\bm{w}_{t-1})$, where $f_{i}(\bm{w}):=\frac{\bm{w}\odot(\mathbf{1}+\bm{R}_{(t-2)n+i})}{1+\bm{w}^{\top}\bm{R}_{(t-2)n+i}}$.
+This method is thus called the cost-aware portfolio estimator (CAPE) in high dimensions.
+This paper proposes a new estimator targeting $\bm{w}^{*}_{1}$ and $\bm{\delta}^{*}_{t}$ under the high-dimensional model setting.
 Figure 1: First portfolio construction and reallocation diagram for multi-periods portfolios selection problems
-We assume the true weight vector $\bm{w}_{1}^{*}$ and weight difference vector $\bm{\delta}^{*}_{t}$ are sparse.
 Denote the support sets of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ as $\mathcal{A}_{1}=\left\{j:w^{*}_{1,j}\neq 0\right\}$ and $\mathcal{A}_{t}=\left\{j:\delta^{*}_{t,j}\neq 0\right\},\ t\geq 2$, respectively.
-And the sparsity assumption means the cardinality $s_{t}=|\mathcal{A}_{t}|\ll p$ for $t=1,2,\ldots,m$.
+And we assume the cardinality $s_{t}=|\mathcal{A}_{t}|\leq s_{0}$, for $t=1,2,\ldots,m$.
 We believe this assumption is empirically grounded as many active strategy investors only choose a relatively small number (compared to the vast universe) of assets considering the transaction and management fees.
-In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to obtain a sparse estimator of $\bm{w}_{1}^{*}$:
-$\widehat{\bm{w}}_{1}:$ $=\mathop{\arg\min}_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})+P_{\lambda}(\bm{w}),$ (2.3)
-where $\widehat{\bm{\Sigma}}_{1}$ and $\widehat{\bm{\mu}}_{1}$ are based on samples $[\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}]^{\top}\in\mathbb{R}^{n\times p}$, $d=n+1$, and $P_{\lambda}(\bm{w})=\sum_{j=1}^{p}P_{\lambda}\left(\bm{w}_{j}\right)$ is the penalty function.
-The oracle (when the true support set $\mathcal{A}_{1}$ is known) portfolio construction weight estimator is given as
-$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle}}=\left(\widehat{\bm{\bm{w}}}_{\mathcal{A}_{1}}^{\text{oracle}},\mathbf{0}\right)=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}).$
-We assume the oracle solution is unique, namely,
-$\nabla_{j}\mathcal{L}_{1,t}\left(\widehat{\bm{\bm{w}}}^{\text{oracle}}_{1}\right)+h^{\text{oracle}}=0\quad\forall j\in\mathcal{A}_{1},$ (2.4)
-where $h^{\text{oracle}}=\frac{2-\gamma\mathbf{1}_{\mathcal{A}_{1}}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{\mu}_{\mathcal{A}_{1}}}{\mathbf{1}_{\mathcal{A}_{1}}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\mathbf{1}_{\mathcal{A}_{1}}}$ is the Lagrange parameter and $\mathcal{L}_{n,1}(\bm{w})=\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})$.
-The above discussion pertains to the construction of oracle estimator $\widehat{\bm{w}}_{1}^{\text{oracle}}$ during the first portfolio construction stage.
-For the reallocation stage, by replacing $\bm{w}$ with $\bm{\delta}$ and substituting the subscript $1$ with $t$, the explanation extends naturally to the construction of oracle estimator $\widehat{\bm{\delta}}_{t}^{\text{oracle}}$ for the reallocation stage.
-A popular choice of $P_{\lambda}(\cdot)$ is the Lasso (Tibshirani, 1996).
-However, the $\ell_{1}$ penalty is biased, requiring a strong Irrepresentable Condition to be selection consistent (Zhao and Yu, 2006).
-Nonconvex penalization method may outperform the $\ell_{1}$ method in terms of estimation accuracy and selection consistency.
-The nonconvex penalty function $P_{\lambda}:\mathbb{R}\rightarrow\mathbb{R}$ satisfies the following conditions:
+To address the computational intractability of the $\ell_{0}$-constrained problems in (2.4) and (2.5), we approximate the $\ell_{0}$ constraint with a general sparsity-inducing penalty $p_{\lambda}(\cdot)$, yielding the following penalized formulations.
+At the initial portfolio construction stage ($t=1$):
+$\widehat{\bm{w}}_{1}^{\mathrm{pen}}=\mathop{\arg\min}_{\bm{w}\in\mathbb{R}^{p}}\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\,\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w})+\sum_{j=1}^{p}p_{\lambda}(|w_{j}|)\quad\text{s.t.}\quad\bm{w}^{\top}\mathbf{1}=1.$ (2.6)
+At the reallocation stage ($t\geq 2$):
+$\widehat{\bm{\delta}}_{t}^{\mathrm{pen}}=\mathop{\arg\min}_{\bm{\delta}\in\mathbb{R}^{p}}\bm{\delta}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}-\gamma\,\bm{\delta}^{\top}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta})+\sum_{j=1}^{p}p_{\lambda}(|\delta_{j}|)\quad\text{s.t.}\quad\bm{\delta}^{\top}\mathbf{1}=0.$ (2.7)
+Here, $p_{\lambda}(\cdot)$ denotes the sparsity-inducing penalty, and $\widehat{\bm{\mu}}_{t}$ and $\widehat{\bm{\Sigma}}_{t}$ are the estimators of the mean vector $\bm{\mu}_{t}$ and covariance matrix $\bm{\Sigma}_{t}$, respectively, for $t=1,2,\dots,m$.
+Two popular choices of $p_{\lambda}(\cdot)$ are the $\ell_{1}$ penalty (i.e., Lasso), and the Smoothly Clipped Absolute Deviation (SCAD) penalty.
+While the Lasso penalty is convex and yields efficient computation, it introduces bias and requires the strong Irrepresentable Condition to achieve variable selection consistency (Zhao and Yu, 2006).
+In contrast, nonconvex penalties like SCAD can reduce the estimation bias and achieve better selection consistency under weaker conditions.
+The SCAD penalty (Fan and Li, 2001) satisfies the following regularity conditions:
   • (i)
-    $P_{\lambda}(t)$ is increasing and concave in $t\in[0,\infty)$ with $P_{\lambda}(0)=0$;
+    $P_{\lambda}(t)$ is increasing and concave on $t\in[0,\infty)$ with $P_{\lambda}(0)=0$;
   • (ii)
-    $P_{\lambda}(t)$ is differentiable in $t\in(0,\infty)$ with $P_{\lambda}^{\prime}(0):=P_{\lambda}^{\prime}(0+)\geq a_{1}\lambda$;
+    $P_{\lambda}$ is differentiable on $t\in(0,\infty)$ with $P_{\lambda}^{\prime}(0+)\geq a_{1}\lambda$;
   • (iii)
-    $P_{\lambda}^{\prime}(t)\geq a_{1}\lambda$ for $t\in\left(0,a_{2}\lambda\right]$;
+    $P_{\lambda}^{\prime}(t)\geq a_{1}\lambda$ for $t\in(0,a_{2}\lambda]$;
   • (iv)
-    $P_{\lambda}^{\prime}(t)=0$ for $t\in[a\lambda,\infty)$ with the pre-specified constant $a>a_{2}$, where $a_{1}$ and $a_{2}$ are two fixed positive constants.
-The mild conditions on $P_{\lambda}(\cdot)$ are typically met by a wide range of regularizers, such as SCAD (Fan and Li, 2001) and MCP (Zhang, 2010).
-It is easy to verify that $a_{1}=a_{2}=1$ for the SCAD, and its derivative is defined as
-$P^{{}^{\prime}}_{\lambda}(|\beta|)=\lambda\left\{I(|\beta|\leq\lambda)+\frac{(a\lambda-|\beta|)_{+}}{(a-1)\lambda}I(|\beta|>\lambda)\right\}$
-for some $a>2$, and $P^{{}^{\prime}}_{\lambda}(0+)=\lambda$.
-However, the nonconvex nature of the penalty term presents computational challenges due to the potential multiple local optima.
-To address this issue, Zou and Li (2008) proposed a feasible solution using LLA.
-This approach transforms the nonconvex penalized model into a sequence of weighted $\ell_{1}$ models, thereby leveraging the computational efficiency of the $\ell_{1}$ method.
-Zou and Li (2008) suggested using the properly initialized one-step LLA estimator instead of the computationally intensive fully converged estimator.
-It has been demonstrated that the one-step LLA estimator enjoys the strong oracle property when $\bm{X}^{\top}\bm{X}/n\rightarrow\mathbf{M}$ for some positive definition matrix $\mathbf{M}$.
-Subsequently, Fan et al.
-(2014) extended this study to high-dimensional settings and showed that initializing the LLA algorithm with a Lasso solution leads to an oracle solution with high probability.
-The computation algorithm for first portfolio construction is summarized in Algorithm 1.
-For the computation algorithm at reallocation stages, it suffices to replace $\bm{w}$ with $\bm{\delta}$ and substitute the subscript $1$ with $t$.
+    $P_{\lambda}^{\prime}(t)=0$ for $t\in[a\lambda,\infty)$ with some constant $a>a_{2}$,
+where $a_{1}$ and $a_{2}$ are fixed positive constants.
+The SCAD penalty has $a_{1}=a_{2}=1$ and its derivative is given by
+$P^{\prime}_{\lambda}(|\tau|)=\lambda\left\{I(|\tau|\leq\lambda)+\frac{(a\lambda-|\tau|)_{+}}{(a-1)\lambda}I(|\tau|>\lambda)\right\}$
+for some $a>2$, and $P^{\prime}_{\lambda}(0+)=\lambda$.
+Building on the general penalized formulations in (2.6) and (2.7), and leveraging the discussion of commonly used penalties such as Lasso and SCAD, we develop a novel cost-aware optimization framework specifically designed for high-dimensional portfolio selection.
+Specifically, we define the Cost-Aware Portfolio Estimator under SCAD Penalty (CAPE-S) at the first portfolio construction stage as follows:
+$\widehat{\bm{w}}_{1}=\mathop{\arg\min}_{\bm{w}_{1}\in\mathbb{R}^{p}}\;\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}_{1})+P_{\lambda}(\bm{w}_{1})\quad\text{s.t.}\quad\bm{w}_{1}^{\top}\mathbf{1}=1,$ (2.8)
+where $P_{\lambda}(\bm{w}_{1})=\sum_{j=1}^{p}P_{\lambda}(\bm{w}_{t,j})$ is SCAD penalty.
+This formulation simultaneously accounts for portfolio risk, expected return, transaction cost, and sparsity, making it particularly suitable for high-dimensional financial settings.
+To facilitate theoretical analysis, we consider the oracle estimator which assumes knowledge of the true support set $\mathcal{A}_{1}$.
+The oracle portfolio weight estimator is defined by restricting the weights outside $\mathcal{A}_{1}$ to zero:
+$\widehat{\bm{w}}_{1}^{\mathrm{oracle}}=\mathop{\arg\min}_{\bm{w}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\,\bm{w}^{\top}_{1}\mathbf{1}=1}\bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}_{1}).$
+We assume the oracle solution satisfies the first-order optimality conditions:
+$\nabla_{j}\mathcal{L}_{n,1}(\widehat{\bm{w}}_{1}^{\mathrm{oracle}})+h^{\mathrm{oracle}}=0,\quad\forall j\in\mathcal{A}_{1},$
+where the Lagrange multiplier
+$h^{\mathrm{oracle}}=\frac{2-\gamma\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\widehat{\bm{\mu}}_{\mathcal{A}_{1}}+\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{g}_{\mathcal{A}_{1}}}{\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{1}},$
+$\mathbf{g}:=\partial\mathbf{C}_{1}(\bm{w}_{1})/\partial\bm{w}_{1}\ \text{is the vector of sub-derivatives of}\ \mathbf{C}_{1}(\bm{w}_{1})$
+and
+$\mathcal{L}_{n,1}(\bm{w})=\bm{w}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}-\gamma\bm{w}^{\top}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}).$
+The above describes the oracle estimator during the first initial portfolio construction stage.
+Similarly, for the reallocation stage with $t\geq 2$, the Cost-Aware Portfolio Estimator under SCAD Penalty (CAPE-S) are defined as
+$\widehat{\bm{\delta}}_{t}=\mathop{\arg\min}_{\bm{\delta}_{t}\in\mathbb{R}^{p}}\;\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta}_{t})+P_{\lambda}(\bm{\delta}_{t})\quad\text{s.t.}\quad\bm{\delta}^{\top}_{t}\mathbf{1}=0,$
+where $P_{\lambda}(\bm{\delta}_{t})=\sum_{j=1}^{p}P_{\lambda}(\bm{\delta}_{t,j})$ is SCAD penalty and the weight difference vector $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, for $t\geq 2$, and $\widehat{\bm{\mu}}_{t}$ and $\widehat{\bm{\Sigma}}_{t}$ denote estimators of the true mean vector $\bm{\mu}_{t}$ and covariance matrix $\bm{\Sigma}_{t}$, respectively.
+The corresponding oracle estimator is
+$\widehat{\bm{\delta}}_{t}^{\mathrm{oracle}}=\mathop{\arg\min}_{\bm{\delta}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\,\bm{\delta}^{\top}_{t}\mathbf{1}=0}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta}_{t}),$
+which satisfies
+$\nabla_{j}\mathcal{L}_{n,t}(\widehat{\bm{\delta}}_{t}^{\mathrm{oracle}})+h_{t}^{\mathrm{oracle}}=0,\quad\forall j\in\mathcal{A}_{t},$
+where
+$h_{t}^{\mathrm{oracle}}=\frac{2-\gamma\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\widehat{\bm{\mu}}_{\mathcal{A}_{t}}+\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{g}_{\mathcal{A}_{t}}}{\mathbf{1}^{\top}\widehat{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\mathbf{1}},$
+$\mathbf{g}:=\partial\mathbf{C}_{t}(\bm{\delta}_{t})/\partial\bm{\delta}_{t}\ \text{is the vector of sub-derivatives of}\ \mathbf{C}_{t}(\bm{\delta}_{t}),$
+and
+$\mathcal{L}_{n,t}(\bm{\delta}_{t})=\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}_{t}\widehat{\bm{\mu}}_{t}+\mathbf{C}_{t}(\bm{\delta}_{t}).$
+However, the nonconvexity of SCAD poses significant computational challenges due to the presence of multiple local minima.
+To overcome this issue, Zou and Li (2008) proposed the local linear approximation (LLA) algorithm, which iteratively approximates the nonconvex problem by a sequence of weighted $\ell_{1}$-penalized problems, thereby inheriting the computational advantages of convex optimization.
+Rather than pursuing full convergence, they advocated for a properly initialized one-step LLA estimator, which is computationally efficient and shown to possess the strong oracle property under the condition that $\bm{X}^{\top}\bm{X}/n\to\mathbf{M}$ for some positive definite matrix $\mathbf{M}$.
+This framework was later extended by Fan et al.
+(2014) to high-dimensional settings, demonstrating that initializing the LLA algorithm with a Lasso solution yields an oracle estimator with high probability.
+The algorithmic procedures for the first portfolio construction and subsequent reallocation stages are presented in Algorithm 1 and Algorithm 2, respectively.
+Here, $\widehat{\bm{w}}_{1}^{\text{initial}}$ and $\widehat{\bm{\delta}}_{t}^{\text{initial}}$ denote initial estimators of $\bm{w}^{*}_{1}$ and $\bm{\delta}^{*}_{t}$, respectively, satisfying certain regularity conditions.
+The specific definitions of these initial estimators, tailored to different transaction cost structures, are provided in Sections 2.3 and 2.4.
 Algorithm 1 Local Linear Approximation (LLA) Algorithm for First Portfolio Construction
 1:Initialize $\widehat{\bm{w}}_{1}^{(0)}=\widehat{\bm{w}}_{1}^{\text{initial}}$ and compute the adaptive weight:
-$\widehat{\bm{\theta}}^{(0)}=\left(\widehat{\theta}_{1}^{(0)},\ldots,\widehat{\theta}_{p}^{(0)}\right)^{\prime}=\left(P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,1}^{(0)}\right|\right),\ldots,P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,p}^{(0)}\right|\right)\right)^{\prime}$
-2:for $k=1,2,...$ do repeat the LLA iteration till convergence
-3:     Obtain $\widehat{\bm{w}}_{t}^{(k)}$ by solving the following optimization problem
-$\widehat{\bm{w}}_{1}^{(k)}=\min_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\mathcal{L}_{n,1}(\bm{\bm{w}})+\sum_{j}\widehat{\theta}_{j}^{(k-1)}\cdot\left|w_{j}\right|$
-4:     Update the adaptive weight vector $\widehat{\bm{\bm{\theta}}}^{(k)}$ with $\widehat{\theta}_{j}^{(k)}=P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,j}^{(k)}\right|\right)$
+$\widehat{\bm{\theta}}^{(0)}=\left(\widehat{\theta}_{1}^{(0)},\ldots,\widehat{\theta}_{p}^{(0)}\right)^{\top}=\left(P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,1}^{(0)}\right|\right),\ldots,P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,p}^{(0)}\right|\right)\right)^{\top}$
+2:for $l=1,2,$... do repeat the LLA iteration till convergence
+3:  Obtain $\widehat{\bm{w}}_{t}^{(l)}$ by solving the following optimization problem
+$\widehat{\bm{w}}_{1}^{(l)}=\min_{\bm{w}\in\mathbf{R}^{p},\bm{w}^{\top}\mathbf{1}=1}\mathcal{L}_{n,1}(\bm{\bm{w}})+\sum_{j}\widehat{\theta}_{j}^{(l-1)}\cdot\left|w_{j}\right|$
+4:  Update the adaptive weight vector $\widehat{\bm{\bm{\theta}}}^{(l)}$ with $\widehat{\theta}_{j}^{(l)}=P_{\lambda}^{\prime}\left(\left|\widehat{w}_{1,j}^{(l)}\right|\right)$
 5:end for
 6:Output: The portfolio allocation $\widehat{\bm{w}}_{t}^{\text{LLA}}$
-For $t=1,2,\ldots,m$, we assume that the estimators for mean and covariance matrix satisfy the following element-wise max norm deviation condition for some absolute constants $c_{1},c_{2},M_{1},M_{2}$:
-$\operatorname{Pr}\left(\|\widehat{\bm{\mu}}_{t}-\bm{\mu}_{t}\|_{\infty}\leq M_{1}\sqrt{\log p/n}\right)\leq 1-O(p^{-c_{1}}),$ (2.5)
-$\operatorname{Pr}\left(\|\widehat{\bm{\Sigma}}_{t}-\bm{\Sigma}_{t}\|_{\max}\leq M_{2}\sqrt{\log p/n}\right)\leq 1-O(p^{-c_{2}})$
-When the returns $\bm{R}_{d-n},\bm{R}_{d-n+1},\ldots,\bm{R}_{d-1}$ are independent and identically distributed sub-Gaussian random vectors with mean $\bm{\mu}$ and covariance $\bm{\Sigma}$, condition 2.5 is satisfied by sample mean and covariance matrix (see Lemma S.1 in the supplementary materials).
-Moreover, in high-dimensional settings, theoretically refined estimators, such as linear and nonlinear shrinkage estimators, are expected to better satisfy the assumptions outlined above.
-Notice that we can relax condition 2.5 and allow dependent return data, including strong-mixing ($\alpha$-mixing), $m$-dependent sequence, and physical (or functional) dependence measure for casual time series (Chang et al., 2024).
-Nonetheless, we do not provide the full theory result of CAPE under dependent return data here for simplicity.
-We further assume the estimator of covariance matrix $\widehat{\bm{\Sigma}}_{t}$ satisfies the following restricted strong convexity (RSC) condition:
-$\forall\mathbf{\Delta}\in\mathbb{R}^{p},\mathbf{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\mathbf{\Delta}\geq v\|\bm{\Delta}\|_{2}^{2}-\tau\sqrt{\frac{\log p}{n}}\|\bm{\Delta}\|_{1}~{}~{}\text{where}~{}~{}v>0,\tau\geq 0.$ (2.6)
-The sample covariance matrix is typically semi-positive definite and $\forall\bm{\Delta}\in\mathbb{R}^{p},\bm{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\Delta}\geq 0$, so the RSC condition hold trivially for $\frac{\|\bm{\Delta}\|_{1}}{\|\bm{\Delta}\|_{2}^{2}}\geq\frac{v}{\tau}\sqrt{\frac{n}{\log p}}$.
-We only require the condition to hold over a cone of the form $\frac{\|\bm{\Delta}\|_{1}}{\|\bm{\Delta}\|_{2}^{2}}\leq\frac{v}{\tau}\sqrt{\frac{n}{\log p}}$.
-The RSC condition is widely adopted in high-dimensional estimation and is satisfied with high probability if the return vector is i.i.d.
-sub-Gaussian random variables (Loh and Wainwright, 2015).
+Algorithm 2 Local Linear Approximation (LLA) Algorithm for Reallocation
+1:Initialize $\widehat{\bm{\delta}}_{t}^{(0)}=\widehat{\bm{\delta}}_{t}^{\text{initial}}$ and compute the adaptive weight:
+$\widehat{\bm{\theta}}^{(0)}=\left(\widehat{\theta}_{1}^{(0)},\ldots,\widehat{\theta}_{p}^{(0)}\right)^{\top}=\left(P_{\lambda}^{\prime}\left(\left|\widehat{\delta}_{1,1}^{(0)}\right|\right),\ldots,P_{\lambda}^{\prime}\left(\left|\widehat{\delta}_{1,p}^{(0)}\right|\right)\right)^{\top}$
+2:for $l=1,2,$... do repeat the LLA iteration till convergence
+3:  Obtain $\widehat{\bm{\delta}}_{t}^{(l)}$ by solving the following optimization problem
+$\widehat{\bm{\delta}}_{t}^{(l)}=\min_{\bm{\delta}\in\mathbf{R}^{p},\bm{\delta}^{\top}\mathbf{1}=0}\mathcal{L}_{n,t}(\bm{\bm{\delta}})+\sum_{j}\widehat{\theta}_{j}^{(l-1)}\cdot\left|\delta_{j}\right|$
+4:  Update the adaptive weight vector $\widehat{\bm{\bm{\theta}}}^{(l)}$ with $\widehat{\theta}_{j}^{(l)}=P_{\lambda}^{\prime}\left(\left|\widehat{\delta}_{1,j}^{(l)}\right|\right)$
+5:end for
+6:Output: The portfolio allocation $\widehat{\bm{\delta}}_{t}^{\text{LLA}}$
 To establish the non-asymptotic error bounds, we impose the following assumptions:
   • (A1)
-    The estimators $\widehat{\bm{\mu}}_{t}$ and $\widehat{\bm{\Sigma}}_{t}$ the estimators satisfy the element-wise max norm deviation condition (2.5) and the RSC condition (2.6);
+    There exist constants $C_{1},C_{2}>0$, for $t=1,2,\ldots,m$, such that
+    $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{r}\|_{\infty}<C_{1},\quad\forall\bm{r}\in\{-1,0,1\}^{s_{t}},$
+[…]
+    $0<\lambda_{\min}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})\leq\lambda_{\max}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})<\infty.$
   • (A2)
-    There exists $C_{1},C_{2}>0$, s.t.
-$\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}\|_{1}<C_{1}$, $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{r}\|_{\infty}<C_{2},\ \forall\bm{r}\in\{-1,0,1\}^{s_{t}}$;
+    For $t=1,2,\ldots,m$ and some absolute constants $c_{1},c_{2},M_{1},M_{2}$,
+    $\operatorname{Pr}\left(\|\widehat{\bm{\mu}}_{t}-\bm{\mu}_{t}\|_{\infty}\leq M_{1}\sqrt{\log p/n}\right)$ $\geq 1-O(p^{-c_{1}}),$ (2.9)
+    $\operatorname{Pr}\left(\|\widehat{\bm{\Sigma}}_{t}-\bm{\Sigma}_{t}\|_{\max}\leq M_{2}\sqrt{\log p/n}\right)$ $\geq 1-O(p^{-c_{2}}).$
   • (A3)
-    The covariance matrix satisfies $0<\lambda_{\min}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})\leq\lambda_{\max}(\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}})<\infty$ and there exists $C_{3}>0$, such that $|||\bm{\Sigma}_{\mathcal{A}^{c}_{t},\mathcal{A}_{t}}|||_{\infty}<C_{3}$;
-  • (A4)
-    Suppose the theoretical minimizer $\bm{w}_{1}^{*}$ is sparse with cardinality $s_{1}=\|\bm{w}^{*}_{1}\|_{0}$ and the minimal signal strength of $\bm{w}_{1}^{*}$ satisfies that $\|\bm{\bm{w}}_{\mathcal{A}_{1}}^{*}\|_{\min}>(a+1)\lambda$.
-Assumptions (A1)–(A3) hold for both the first portfolio construction stage and the reallocation stage, implying that $t=1,2,\ldots,m$.
-As demonstrated by Liu et al.
-(2024), the sample mean and covariance matrix satisfy the condition (2.5) in Assumption (A1), if the returns $\bm{R}_{d-n},\ldots,\bm{R}_{d-1}$ are i.i.d.
-sub-Gaussian random vectors with mean $\bm{\mu}$ and covariance matrix $\bm{\Sigma}$, and $\log p<n$.
-To improve the estimation accuracy, both linear and nonlinear shrinkage estimator employed in this study are assumed to satisfy condition (2.5).
-In Assumption (A2), for simplicity, we denote the submatrix of $\bm{\Sigma}_{t}$ consisting of the rows and columns indexed by $\mathcal{A}_{t}$ as $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$, and similarly, the subvector of $\bm{\mu}_{t}$ containing the elements indexed by $\mathcal{A}_{t}$ is represented as $\bm{\mu}_{\mathcal{A}_{t}}$.
-This notation will be adopted consistently in the subsequent discussion.
-Assumption (A2) is introduced to ensure that the $\ell_{1}$-norm of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ remain bounded, thereby preventing divergence.
-Similar assumptions can be found in Cai et al.
-(2022) and Liu et al.
+    The estimated covariance matrix $\widehat{\bm{\Sigma}}_{t}$ satisfies the restricted strong convexity (RSC) condition:
+    $\forall\bm{\Delta}\in\mathbb{R}^{p},\quad\bm{\Delta}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\Delta}\geq v\|\bm{\Delta}\|_{2}^{2}-\tau\sqrt{\frac{\log p}{n}}\|\bm{\Delta}\|_{1},$ (2.10)
+    where $v>0$, $\tau\geq 0$, and $t=1,2,\ldots,m$.
+Remark 1.
+Assumption (A1) guarantees three key properties.
+The bound on $\|\bm{\Sigma}^{-1}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{r}\|_{\infty}$ ensures that the $\ell_{1}$-norms of $\bm{w}_{1}^{*}$ and $\bm{\delta}_{t}^{*}$ are controlled, preventing divergence in the estimation process.
+This condition aligns with assumptions in Cai et al.
 (2024).
-And the assumption constraining the $\ell_{1}$-norm of the parameter to prevent divergence also appears in Loh and Wainwright (2015).
-Furthermore, Assumption (A3) is less restrictive than the Irrepresentable Condition required for the Lasso, which restricts the strength of interactions between variables in $\mathcal{A}_{t}$ and $\mathcal{A}^{c}_{t}$.
-It ensures that the influence of variables outside $\mathcal{A}_{t}$ are not excessively large, preserving sparsity or stability in the covariance structure.
-Finally, the sparsity assumption in Assumption (A4) is reasonable and has also been employed in Du et al.
+The constraint on $|\bm{\Sigma}_{\mathcal{A}^{c}_{t},\mathcal{A}_{t}}|||_{\infty}$, while weaker than the irrepresentable condition required for the Lasso, limits the correlation between active and inactive variables, promoting both sparsity and stability.
+Finally, the eigenvalue condition guarantees that $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$ is well-conditioned, which facilitates stable matrix inversion and avoids ill-posedness when estimating coefficients restricted to the active set $\mathcal{A}_{t}$.
+Assumption (A2) holds when the sample mean vector and sample covariance matrix are used as estimators, provided that the returns $\bm{R}_{(t-1)n+1},\ldots,\bm{R}_{tn}$ are i.i.d.
+sub-Gaussian vectors with mean $\bm{\mu}_{t}$ and covariance matrix $\bm{\Sigma}_{t}$, and that $\log p<n$, as shown in Liu et al.
 (2024).
-Assumption (A4) also imposes the minimal signal strength condition, which is required for selection consistency, and the similar condition can be found in Fan et al.
-(2014), Loh and Wainwright (2017).
-2.2 CAPE with Quadratic Transaction Cost
+Furthermore, we show that the linear shrinkage estimator (Ledoit and Wolf, 2020) for the population covariance matrix also satisfies (A2) under the same sub-Gaussianity assumption and mild structural conditions on $\bm{\Sigma}_{t}$ in the supplementary material.
+The linear shrinkage estimator has more favorable properties and, in high-dimensional settings (e.g., $p$ comparable to or exceeding $n$), provides well-conditioned covariance estimates that make the classical Markowitz mean–variance model (2.1) and its transaction-cost extension (2.3) feasible and numerically stable.
+Moreover, the bounds in (2.9) can be extended to settings with dependent observations, including strong-mixing ($\alpha$-mixing) processes, $m$-dependence, and time series exhibiting physical (or functional) dependence (Chang et al., 2024).
+For simplicity, we focus on the i.i.d.
+case in this paper and defer the dependent setting to future work.
+Assumption (A3) is a standard requirement in high-dimensional estimation.
+Loh and Wainwright (2015) established that the sample covariance matrix satisfies the RSC condition with high probability under sub-Gaussian assumptions.
+We further show that both linear and nonlinear shrinkage estimators (Ledoit and Wolf, 2020) satisfy (A3), as detailed in the supplementary material.
+2.3 CAPE-S with Quadratic Transaction Cost
 2.3.1 First Portfolio Construction
@@
 $\mathbf{C}_{1}(\bm{w})=(\bm{\beta}\odot\bm{w})^{\top}\bm{w}$
-where $\bm{\beta}=(\beta_{1},...,\beta_{p})^{\top},\beta_{j}>0$ is the cost parameter, which means the unit transaction cost of the risky asset $j$ is a linear function of the trading size with slope $\beta_{j}$, for $j=1,...,p$.
+where $\bm{\beta}=(\beta_{1},...,\beta_{p})^{\top},\beta_{j}>0$ is the cost parameter, which means the unit transaction cost of the risky asset $j$ is a linear function of the trading size with slope $\beta_{j}$, for j=1,,pj=1,...,p.
 The problem (2.4) then becomes
-$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1},\quad\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$
+$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1},\quad\text{ s.t.
+}\bm{w}_{1}^{\top}\mathbf{1}=1,\ \|\bm{w}_{1}\|_{0}\leq s_{0}.$ (2.11)
 This is a constrained convex optimization problem.
-By Lagrange multiplier method and the definition of $\bm{w}^{*}$ and $\mathcal{A}_{t}$, we have the explicit solution $\bm{w}_{1}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{1}^{-1}\left(\gamma\bm{\mu}_{1}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{1}^{-1}\bm{\mu}_{1}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{1}^{-1}\mathbf{1}}\mathbf{1}\right)$, and $\bm{w}^{*}_{\mathcal{A}_{1}^{c}}=\mathbf{0}$, by sparsity assumption, where $\tilde{\bm{\Sigma}}_{1}=\bm{\Sigma}_{1}+\mathbb{B}$, $\mathbb{B}=\text{diag}(\beta_{1},...,\beta_{p})$.
-As discussed earlier, we use the LLA algorithm for SCAD, and $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ is the initial and Lasso solution.
-In the following lemma, we show the consistency of the initial Lasso weight estimator.
+By applying the method of Lagrange multipliers and using the definitions of $\bm{w}^{*}_{1}$ and the active set $\mathcal{A}_{1}$, we obtain the explicit solution
+$\bm{w}_{\mathcal{A}_{1}}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\left(\gamma\bm{\mu}_{\mathcal{A}_{1}}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\bm{\mu}_{\mathcal{A}_{1}}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}^{-1}\mathbf{1}}\mathbf{1}\right)\ \text{and}\ \bm{w}_{\mathcal{A}^{c}_{1}}^{*}=\mathbf{0}.$
+Here, $\tilde{\bm{\Sigma}}_{1}=\bm{\Sigma}_{1}+\mathbb{B}$, with $\mathbb{B}=\mathrm{diag}(\beta_{1},\ldots,\beta_{p})$.
+The explicit solution above is derived using the method of Lagrange multipliers and the sparsity structure of the optimizer.
+For a detailed derivation, we refer the reader to the proof of Theorem 1 in the supplementary material.
+As discussed earlier, we employ the LLA algorithm to solve CAPE-S and use the CAPE-L $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ as the initial estimate required by the LLA algorithm.
+The following lemma establishes the consistency of the CAPE-L.
 Lemma 1.
-For the first portfolio construction period, suppose that Assumptions (A1)-(A3) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$, where $C_{2}$ is defined in Assumption (A2).
+For the first portfolio construction period, suppose that Assumptions (A1)-(A3) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$, where $C_{1}$ is defined in Assumption (A1).
 Let $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ be the unique optimum of the program
-$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}=1}\bm{w}_{1}^{\top}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1}+\lambda_{Lasso}\|\bm{w}\|_{1}$ (2.7)
-with $\lambda_{Lasso}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
-$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}\sqrt{\frac{\log p}{n}}\right),\quad\quad\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{\frac{s_{1}\log p}{n}}\right).$
+$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}_{1}=1}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1}+\lambda_{\text{Lasso},1}\|\bm{w}_{1}\|_{1}$ (2.12)
+with $\lambda_{\text{Lasso},1}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
+$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{\frac{s_{1}\log p}{n}}\right).$
+In Lemma 1, the estimator $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\beta}$ defined in Equation (2.12) essentially applies the $\ell_{1}$ penalty to the penalty term introduced in Equation (2.6).
+Following the naming convention used in Equation (2.8), this class of estimators can be referred to as the Cost-Aware Portfolio Estimator under Lasso Penalty (CAPE-L) at the first portfolio construction stage.
 Lemma 1 says that the initial Lasso estimator yields an estimate that is consistent in $\ell_{2}$ norm if $\sqrt{s_{1}\log p/n}\to 0$ as $n,p\to\infty$.
@@
 For the first portfolio construction, the oracle portfolio construction weight estimator with quadratic transaction cost is given as
-$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{\mathcal{A}_{1}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{w}_{\mathcal{A}_{1}}-\gamma\bm{w}^{\top}_{\mathcal{A}_{1}}\bm{\mu}_{1}+(\bm{\beta}_{{\mathcal{A}_{1}}}\odot\bm{w}_{{\mathcal{A}_{1}}})^{\top}\bm{w}_{\mathcal{A}_{1}}.$ (2.8)
-The following theorem states that we can find the oracle estimator after two iterations by implementing the LLA algorithm, and it also provides the convergence rate in $\ell_{\infty}$-norm.
+$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}_{1}\mathbf{1}=1}\bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+(\bm{\beta}\odot\bm{w}_{1})^{\top}\bm{w}_{1}.$ (2.13)
+The following theorem states that the oracle estimator can be obtained after two iterations by implementing the LLA algorithm to solve CAPE-S, and it also provides the convergence rate in the $\ell_{\infty}$-norm.
 Theorem 1.
-Suppose that Assumptions (A1)-(A4) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2), we choose $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$.
-Let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$.
-If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$.
+Suppose that Assumptions (A1)–(A3) hold and that $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$, where $C_{1}$ is defined in Assumption (A1).
+Choose $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some sufficiently large constant $M$.
+Let $\widehat{\bm{w}}^{\mathrm{LLA},\beta}_{1}$ denote the solution obtained from the LLA algorithm 1 with the SACD penalty, initialized at $\widehat{\bm{\bm{w}}}_{1}^{\mathrm{Lasso},\beta}$.
+If $s_{1}=O\!\left(\sqrt{\frac{n}{\log p}}\right)$ and $\|\bm{w}^{*}_{1}\|_{\min}>(a+1)\lambda_{1}$, then $\widehat{\bm{w}}^{\mathrm{LLA},\beta}_{1}$ converges to $\widehat{\bm{\bm{w}}}_{1}^{\mathrm{oracle},\beta}$ within two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for all $k\in[s_{1},p]$ and some constants $c_{1},c_{2}>0$.
 Furthermore,
-$\|\widehat{\bm{\bm{w}}}_{1}^{\text{LLA},\beta}-\bm{w}_{1}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
-Theorem 1 states that under mild conditions, our estimator agrees with the oracle estimator $\widehat{\bm{w}}_{1}^{\text{oracle},\beta}$ with high probability when $k$ diverges to infinity.
-Hence, we only study the property of non-convex penalty in the following sections.
+$\|\widehat{\bm{\bm{w}}}_{1}^{\mathrm{LLA},\beta}-\bm{w}_{1}^{*}\|_{\infty}=O\!\left(\sqrt{\frac{\log k}{n}}\right).$
+Theorem 1 establishes that, under mild regularity conditions, the proposed CAPE-S estimator attains the same performance as the oracle estimator $\widehat{\bm{w}}_{1}^{\mathrm{oracle},\beta}$ after only two iterations of the LLA algorithm.
+In particular, when the portfolio dimension $k$ diverges, the two estimators coincide with high probability, ensuring that CAPE-S inherits the oracle’s optimality properties in large-scale settings.
+This oracle property provides the theoretical foundation for subsequent results on risk and return estimation.
 Corollary 1.
@@
 2.3.2 Portfolio Reallocation
-At reallocation decision day $d$, the transaction costs $C_{t}(\bm{w})$ are a quadratic function of the weight difference vector $\bm{\delta}_{t}$, defined as $\bm{\delta}_{t}:=\bm{w}_{t}-\bm{w}^{+}_{t-1}$.
+At reallocation decision day $d$, the transaction costs $C_{t}(\bm{w})$ are a quadratic function of the weight difference vector $\bm{\delta}_{t}$, defined as $\bm{\delta}_{t}:=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, for $t\geq 2$.
 And $\bm{\delta}_{t}$ represents the difference between the new allocation $\bm{w}_{t}$ and the market value-adjusted allocation $\bm{w}^{+}_{t-1}$.
 The form of transaction costs $C_{t}(\bm{w})$ is thus given by
-$\mathbf{C}_{t}(\bm{w})=[\bm{\beta}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})]^{\top}(\bm{w}_{t}-\bm{w}^{+}_{t-1})$
+[…]
 The program (2.5) turns out to be
-$\min_{\bm{w}_{t}}\bm{w}^{\top}_{t}\bm{\Sigma}_{t}\bm{w}_{t}-\gamma\bm{w}^{\top}_{t}\bm{\mu}_{t}+[\bm{\beta}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})]^{\top}(\bm{w}_{t}-\bm{w}^{+}_{t-1}),\quad\text{ s.t.
-}\bm{w}_{t}^{\top}\mathbf{1}=1,$ (2.9)
-With $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, the optimization problem has the equivalent form:
-$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t}+(\bm{\beta}\odot\bm{\delta}_{t})^{\top}\bm{\delta}_{t},\quad\text{ s.t.
-}\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.10)
-Here, we need to modify Assumption (A4) and propose the following revised assumption:
-  • (A4’)
-    Suppose the theoretical minimizer $\bm{\delta}_{t}^{*}$ is sparse with cardinality $s_{t}=\|\bm{\delta}^{*}_{t}\|_{0}$ and the minimal signal strength of $\bm{\delta}_{t}^{*}$ satisfies that $\|\bm{\bm{\delta}}_{\mathcal{A}_{t}}^{*}\|_{\min}>(a+1)\lambda$;
-Similarly, using the Lagrange multiplier method, we have that $\bm{\delta}_{t}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{t}^{-1}(\gamma\tilde{\bm{\mu}}_{t}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{t}^{-1}\tilde{\bm{\mu}}_{t}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{t}^{-1}\mathbf{1}}\mathbf{1})$ and $\bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0}$, where $\tilde{\bm{\mu}}_{t}=\bm{\mu}_{t}-\frac{2}{\gamma}\bm{\Sigma}_{t}\bm{w}^{+}_{t-1}$.
+$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t}+(\bm{\beta}\odot\bm{\delta}_{t})^{\top}\bm{\delta}_{t},\text{ s.t.
+}\bm{\delta}_{t}^{\top}\mathbf{1}=0,\ \|\bm{\delta}_{t}\|\leq s_{0}.$ (2.14)
+Similarly, using the Lagrange multiplier method, we have that
+$\bm{\delta}_{\mathcal{A}_{t}}^{*}=\frac{1}{2}\tilde{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\left(\gamma\bm{\mu}_{\mathcal{A}_{t}}-2(\bm{\Sigma}_{t}\bm{w}^{+}_{t-1})_{\mathcal{A}_{t}}+\frac{2-\gamma\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\bm{\mu}_{\mathcal{A}_{t}}}{\mathbf{1}^{\top}\tilde{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\mathbf{1}}\mathbf{1}\right)\ \text{and}\ \bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0}.$
 In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to encourage a sparse estimator of $\bm{\delta}_{t}^{*}$
 $\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t.
-}\bm{\delta}_{t}^{\top}\mathbf{1}=0,$ (2.11)
+}\bm{\delta}_{t}^{\top}\mathbf{1}=0,$
 where $\widehat{\bm{\Sigma}}_{t}$ and $\widehat{\bm{\mu}}_{t}$ are the estimators of covariance matrix and mean of the asset returns of the $t$ period, $\widehat{\tilde{\bm{\Sigma}}}_{t}=\widehat{\bm{\Sigma}}_{t}+\mathbb{B}$.
-By incorporating the transaction costs in the mean variance problem and regularizing the eventual rebalancing events, our portfolio achieves better Sharpe ratio and less transaction costs, which is shown in the numerical studies.
+By incorporating the transaction costs in the mean variance problem and regularizing rebalancing events, our portfolio achieves a higher post-cost Sharpe ratio and reduced transaction costs, as demonstrated in numerical studies.
 Define the oracle estimator (with known true support of $\bm{\delta}$) with quadratic transaction cost under portfolio reallocation as
-$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=1}\bm{\delta}^{\top}_{\mathcal{A}_{t}}\widehat{\tilde{\bm{\Sigma}}}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}+2(\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t})_{\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}-\gamma\bm{\delta}^{\top}_{\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}.$ (2.12)
+$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}=$ $\mathop{\arg\min}_{\bm{\bm{\delta}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=0}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}_{t}\widehat{\bm{\mu}}_{t}.$ (2.15)
+where $\widehat{\tilde{\bm{\Sigma}}}_{t}=\widehat{\bm{\Sigma}}_{t}+\mathbb{B}$, with $\mathbb{B}=\mathrm{diag}(\beta_{1},\ldots,\beta_{p})$.
+Lemma 2.
+Suppose that Assumptions (A1)-(A3) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$ with $C_{1}$ defined in Assumption (A1), let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$ be the unique optimum of the program
+$\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+\lambda_{\text{Lasso},t}\|\bm{\delta}_{t}\|_{1}.$ (2.16)
+with $\lambda_{\text{Lasso},t}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
+$\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{\frac{s_{t}\log p}{n}}\right).$
 Theorem 2.
-Suppose that Assumptions (A1)-(A3) and (A4’) hold and $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2), let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$ be the unique optimum of the program $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\tilde{\bm{\Sigma}}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+\lambda_{Lasso,t}\|\bm{\delta}_{t}\|_{1}$.
-Then we have $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)$ and $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{\frac{s_{t}\log p}{n}}\right)$.
-If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, we select $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$ and let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso }}.$
+Suppose that Assumptions (A1)-(A3) hold, and that $\|\bm{\beta}\|_{\infty}\ll 1/C_{1}$, where $C_{1}$ is defined in Assumption (A1).
+Let $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, and assume that $\left\|\bm{\delta}^{*}_{t}\right\|_{\min}>(a+1)\lambda_{t}$.
+Choose $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some sufficiently large constant $M$.
+Let $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ be the optimum of the LLA algorithm 2 with SCAD penalty, initialized at $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\beta}$, as defined in (2.16).
+Then, we have the following results:
   • (i)
-    If $\|\bm{w}^{+}_{t-1}\|_{1}<\infty$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$.
-Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log p/n}\right).$
+    If $\|\bm{w}^{+}_{t-1}\|_{1}<C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}>0$.
+Furthermore,
+    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log p}{n}}\right).$
   • (ii)
-    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq\infty$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$, $\widehat{\bm{\delta}}^{\text{LLA},\beta}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constant $c_{1},c_{2}$.
-Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log k/n}\right)$.
-Theorem 2 implies that in portfolio reallocation, the LLA algorithm initialized at a Lasso solution yields a local optimum of the nonconvex regularized problem that satisfies oracle properties with low $\ell_{\infty}$ error.
-Theorem 2 highlights the impact of different estimation methods employed in the previous $t-1$ stages on the performance of our proposed method in stage $t$.
-Specifically, if regularized estimation methods, such as the proposed CAPE or Lasso, were consistently applied in the previous $t-1$ stages, the resulting $\bm{w}^{+}_{t-1}$ will exhibit a certain level of sparsity.
+    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq C_{0}$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$ and absolute constant $C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\beta}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for all $k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constants $c_{1},c_{2}>0$.
+Furthermore,
+    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\beta}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
+Theorem 2 and Lemma 2 have implied that in portfolio reallocation, the LLA algorithm initialized at CAPE-L yields a local optimum of the nonconvex regularized problem that satisfies oracle properties with low $\ell_{\infty}$ error.
+Theorem 2 highlights the impact of different estimation methods employed in the previous t1t-1 stages on the performance of our proposed method in stage $t$.
+Specifically, if regularized estimation methods, such as the proposed CAPE-S in (2.8) or CAPE-L in (2.12), were consistently applied in the previous t1t-1 stages, the resulting $\bm{w}^{+}_{t-1}$ will exhibit a certain level of sparsity.
 This sparsity enables $\widehat{\bm{\delta}}^{\text{LLA},\beta}_{t}$ to achieve a superior convergence rate in the current stage.
@@
 However, the convergence rate will be $\sqrt{\log p/n}$, which is attributable to the lack of sparsity in $\bm{w}^{+}_{t-1}$.
-Corollary 2 provides detailed conclusions regarding the estimation of $\bm{w}^{*}_{t}$ when the proposed CAPE method is consistently applied across the previous $t-1$ stages.
+Corollary 2 provides detailed conclusions regarding the estimation of $\bm{w}^{*}_{t}$ when the proposed CAPE-S method is consistently applied across the previous t1t-1 stages.
 Corollary 2.
-Define $\max\{s_{1},\ldots,s_{t}\}\leq k_{t}\leq p$ and $\widehat{\bm{w}}_{t}^{\text{LLA},\beta}=\bm{w}_{t-1}^{+}+\widehat{\bm{\delta}}_{t}^{\text{LLA},\beta}$, and we use CAPE to make estimation in the past $t-1$ stage.
+Define $s_{0}\leq k_{t}\leq p$ and $\widehat{\bm{w}}_{t}^{\text{LLA},\beta}=\bm{w}_{t-1}^{+}+\widehat{\bm{\delta}}_{t}^{\text{LLA},\beta}$, and we use CAPE-S to make estimation in the past t1t-1 stage.
 Given the same conditions as in Theorem 1 and Theorem 2, with $k_{t}\sqrt{\log k_{t}/n}=o(1)$, then we have
@@
 Corollary 2 shows that if the sparsity level and the dimension satisfy that $k_{t}\sqrt{\log k_{t}/n}=o(1)$, the in-sample and out-of-sample Shape Ratio estimation errors of reallocation periods vanish when $n$ goes to infinity.
-2.3 CAPE with Proportional Transaction Cost
+2.4 CAPE-S with Proportional Transaction Cost
 2.4.1 First Portfolio Construction
@@
 For the proportional transaction cost, the optimization problem corresponding to the initial period is as follows:
-$\min_{\bm{w}_{1}}\bm{w}^{\top}_{1}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$ (2.13)
+$\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1},\text{ s.t.
+}\bm{w}_{1}^{\top}\mathbf{1}=1,\ \|\bm{w}_{1}\|_{0}\leq s_{0},$
 where $\bm{\alpha}=(\alpha_{1},...,\alpha_{p})^{\top},\alpha_{j}>0$ is the cost parameter, which is similar to $\bm{\beta}$ in Section 2.3.
-The first-order condition of the constrained optimization (2.13) is
-$2\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{\mu}_{1}+\bm{\alpha}\odot\mathbf{g}-h^{*}\mathbf{1}=0,$ (2.14)
-where $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{1}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{1}\right\|_{1}/\partial\bm{w}_{1}$, consisting of elements which are 1 or $-1$ in case $w_{j}>0$ or $w_{j}<0$, respectively, or $\in[-1,1]$ in case $w_{j}=0$, and $h^{*}=\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{1}^{-1}\bm{\mu}_{1}+\mathbf{1}^{\top}\bm{\Sigma}_{1}^{-1}(\bm{\alpha}\odot\mathbf{g})}{\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\mathbf{1}}$ denotes the Lagrangian multiplier.
 By Lagrange multiplier method and the definition of $\bm{w}^{*}_{1}$ and $\mathcal{A}_{1}$, solving for $\bm{w}_{1}^{*}$ yields
-$\bm{w}_{1}^{*}=\frac{1}{2}\bm{\Sigma}^{-1}_{t}\left(\gamma\bm{\mu}_{1}-\bm{\alpha}\odot\mathbf{g}+\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}\bm{\mu}_{1}+\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}(\bm{\alpha}\odot\mathbf{g})}{\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{w}^{*}_{\mathcal{A}^{c}_{1}}=\mathbf{0}.$
-In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to encourage a sparse estimator of $\bm{w}^{*}_{1}$:
-$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+P_{\lambda}(\bm{w}_{1}),\text{ s.t.
-}\bm{w}_{1}^{\top}\mathbf{1}=1.$ (2.15)
-Similar to Lemma 1 in Section 2.2, Lemma 2 shows the consistency of the initial Lasso weight estimator under proportional transaction cost.
+$\bm{w}_{\mathcal{A}_{1}}^{*}=\frac{1}{2}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\left(\gamma\bm{\mu}_{\mathcal{A}_{1}}-\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{1}}+\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{\mu}_{\mathcal{A}_{1}}+\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}(\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{1}})}{\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{w}^{*}_{\mathcal{A}^{c}_{1}}=\mathbf{0}.$
+where $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{1}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{1}\right\|_{1}/\partial\bm{w}_{1}$, consisting of elements which are 1 or 1-1 in case $w_{j}>0$ or $w_{j}<0$, respectively, or $\in[-1,1]$ in case $w_{j}=0$.
+Similar to Lemma 1 in Section 2.3, Lemma 3 shows the consistency of the CAPE-L $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}$ under proportional transaction cost.
 Lemma 3.
 For the initial period, suppose that Assumptions (A1)-(A3) hold, let $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}$ be the unique optimum of the program
-$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}_{1}=1}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+\lambda_{\text{Lasso}}\|\bm{w}_{1}\|_{1},$ (2.16)
-with $\lambda_{\text{Lasso}}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$, then with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}$, we have
-$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}(\alpha+\sqrt{\frac{\log p}{n}})\right),\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{s_{1}}(\alpha+\sqrt{\frac{\log p}{n}})\right),$
-where $\theta$ is the minimal eigenvalue of $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$ and $\alpha=\|\bm{\alpha}\|_{\infty}$.
+$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}_{1}=1}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+\lambda_{\text{Lasso},1}\|\bm{w}_{1}\|_{1},$ (2.17)
+with $\lambda_{\text{Lasso},1}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$, then with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}$, we have
+$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}\|\bm{\alpha}\|_{\infty}+s_{1}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{s_{1}}\|\bm{\alpha}\|_{\infty}+\sqrt{\frac{s_{1}\log p}{n}}\right).$
 Lemma 3 differs from Lemma 1 in Section 2.3 due to the distinct effects of the quadratic and proportional transaction costs.
@@
 Similar to (2.13), at first portfolio construction stage, we define the oracle portfolio construction weight estimator with proportional transaction cost as
-$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\alpha}=\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{\mathcal{A}_{1}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{w}_{\mathcal{A}_{1}}-\gamma\bm{w}^{\top}_{\mathcal{A}_{1}}\bm{\mu}_{1}+\|\bm{\alpha}_{\mathcal{A}_{1}}\odot\bm{w}^{\top}_{\mathcal{A}_{1}}\|_{1}.$ (2.17)
+$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\alpha}=\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}.$ (2.18)
 Theorem 3.
-Under Assumptions (A1)-(A4), we select $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$.
-Let $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{w}}}_{t}^{\text{Lasso },\alpha}$.
-If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$ and $\|\bm{\alpha}\|_{\infty}\leq C\sqrt{\frac{\log p}{n}}$ for some large constant $C$, $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{w}}}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$.
+Under Assumptions (A1)-(A3) and $\|\bm{\alpha}\|_{\infty}\leq C\sqrt{\frac{\log p}{n}}$ for some large constant $C$, we select $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$.
+Let $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm 1 with SCAD penalty, initialized at $\widehat{\bm{\bm{w}}}_{t}^{\text{Lasso },\alpha}$.
+If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$ and $\left\|\bm{w}^{*}_{1}\right\|_{\min}>(a+1)\lambda_{1}$, $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{w}}}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$.
 Furthermore,
@@
 This differs from the assumption in Section 2.3, where $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A1).
-This distinction is also evident in our later sections on simulation studies and real data applications, where $\bm{\beta}=0.15\times\mathbf{1}_{p}$ while $\bm{\alpha}=0.001\times\mathbf{1}_{p}$.
 Hautsch and Voigt (2019) examined the empirical implications of proportional transaction costs on the Sharpe ratio, demonstrating that smaller values of $\alpha$ yield optimal Sharpe ratios.
 Our choice of $\alpha$ aligns with this finding, maintaining the same order of magnitude.
-Building on the insights from Corollary 1, we derive that under the condition $s_{1}\sqrt{s_{1}\log p}=o(1)$, the estimation errors for in-sample and out-of-sample Sharpe ratios during reallocation periods asymptotically vanish as $n$ approaches infinity.
+As in Corollary 1, we can show that under the condition $s_{1}\sqrt{s_{1}\log p}=o(1)$, the estimation errors for in-sample and out-of-sample Sharpe ratios during reallocation periods asymptotically vanish as $n$ approaches infinity.
+Since the conclusion is identical to that of Corollary 1, we omit the formal statement here.
 This result highlights the diminishing impact of estimation uncertainty in high-dimensional settings when the sparsity condition is adequately controlled.
@@
 The program (2.5) turns out to be
-$\min_{\bm{w}_{t}}\bm{w}^{\top}_{t}\bm{\Sigma}_{t}\bm{w}_{t}-\gamma\bm{w}^{\top}_{t}\bm{\mu}_{t-1}+\|\bm{\alpha}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})\|_{1}\text{ s.t. }\bm{w}_{t}^{\top}\mathbf{1}=1,$
-for $t=2,3,...m$, with $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, which has the equivalent form:
-$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1},\text{ s.t.
-}\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.18)
+$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1},\text{ s.t.
+}\bm{\delta}_{t}^{\top}\mathbf{1}=0,\ \|\bm{\delta}_{t}\|_{0}\leq s_{0}.$
 Similarly, solving for $\bm{\delta}_{t}^{*}$ yields
-$\bm{\delta}_{t}^{*}=\frac{1}{2}\bm{\Sigma}_{t}^{-1}\left(\gamma\tilde{\bm{\mu}}_{t}-\bm{\alpha}\odot\mathbf{g}+\frac{2+\gamma\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\bm{\alpha}\odot\mathbf{g}-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\tilde{\bm{\mu}}_{t}}{\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0},$
-where $\tilde{\bm{\mu}}=\bm{\mu}-\frac{2}{\gamma}\bm{\Sigma}_{t}\bm{w}^{+}_{t-1}$ and $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{t}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{t}\right\|_{1}/\partial\bm{w}_{t}$.
-In practice, the feasible form is
-$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\widehat{\bm{w}}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t.
-}\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.19)
+$\bm{\delta}_{\mathcal{A}_{t}}^{*}=\frac{1}{2}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\left(\gamma\bm{\mu}_{\mathcal{A}_{t}}-2(\bm{\Sigma}_{t}\bm{w}^{+}_{t-1})_{\mathcal{A}_{t}}-\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{t}}+\frac{2+\gamma\mathbf{1}^{\top}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{t}}-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\bm{\mu}_{\mathcal{A}_{t}}}{\mathbf{1}^{\top}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\mathbf{1}}\mathbf{1}\right)\ \text{and}\ \bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0},$
+where $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{t}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{t}\right\|_{1}/\partial\bm{w}_{t}$. In practice, the feasible form is
+$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t.
+}\bm{\delta}_{t}^{\top}\mathbf{1}=0.$
 Similar to (2.15), we define the oracle estimator with proportional transaction cost under portfolio reallocation as
-$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=1}\bm{\delta}^{\top}_{\mathcal{A}_{t}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}+2(\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t})_{\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}-\gamma\bm{\delta}^{\top}_{\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}+\|\bm{\alpha}_{\mathcal{A}_{t}}\odot\bm{\delta}_{\mathcal{A}_{t}}\|_{1}.$ (2.20)
+$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}=$ $\mathop{\arg\min}_{\bm{\bm{\delta}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}_{t}^{\top}\mathbf{1}=0}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}_{t}\widehat{\bm{\mu}}_{t}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}.$ (2.19)
+Lemma 4.
+Suppose that Assumptions (A1)–(A3) hold, and that $\|\bm{\alpha}\|_{\infty}=C\sqrt{\frac{\log p}{n}}$ for some large constant $C$.
+Let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}$ denote the unique solution to the following optimization problem:
+$\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}_{t}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+\lambda_{\text{Lasso},t}\|\bm{\delta}_{t}\|_{1}.$ (2.20)
+with $\lambda_{\text{Lasso},t}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
+$\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{\frac{s_{t}\log p}{n}}\right).$
 Theorem 4.
-Suppose that Assumptions (A1)-(A3) and (A4’) hold and $\|\bm{\alpha}\|_{\infty}=C\sqrt{\frac{\log p}{n}}$ for some large constant $C$.
-Let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}$ be the unique optimum of the program $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\widehat{\bm{w}}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+\lambda_{Lasso,t}\|\bm{\delta}_{t}\|_{1}$.
-Then we have $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)$ and $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{s_{t}\frac{\log p}{n}}\right)$.
-If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, we choose $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$ and let $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}.$
+Suppose that Assumptions (A1)–(A3) hold and that $\|\bm{\alpha}\|_{\infty}=C\sqrt{\frac{\log p}{n}}$ for some large constant $C$.
+If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$ and $\|\bm{\delta}_{t}^{*}\|_{\min}>(a+1)\lambda_{t}$, choose $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$.
+Let $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ be the optimum of the LLA algorithm 2 with SCAD penalty, initialized at $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}$ as defined in (2.20).
+Then, the following holds:
   • (i)
-    If $\|\bm{w}^{+}_{t-1}\|_{1}<C_{4}$, $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$.
-Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log p/n}\right).$
+    If $\|\bm{w}^{+}_{t-1}\|_{1}<C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}>0$.
+Furthermore,
+    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log p}{n}}\right).$
   • (ii)
-    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq C_{4}$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$, $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constant $c_{1},c_{2}$.
-Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log k/n}\right)$.
-Remark.
+    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq C_{0}$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$ and absolute constant $C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for all $k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constants $c_{1},c_{2}>0$.
+Furthermore,
+    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
+Remark 2.
 Although the conditions and conclusions of Theorem 4 and Theorem 2 appear to be very similar, the differences arise from the distinct assumptions regarding transaction costs.
@@
 This distinction stems from the differing impacts of transaction costs on the optimization problem introduced in Section 2.2.
-For quadratic transaction costs, it suffices to verify that the modified covariance matrix continues to satisfy our assumptions, as demonstrated in the Lemma LABEL:Property_of_tilde_bSigma provided in the Supplementary Material.
+For quadratic transaction costs, it suffices to verify that the modified covariance matrix continues to satisfy our assumptions, as demonstrated in the Supplementary Material.
 However, for proportional transaction costs, Lemma 3 explicitly illustrates the influence of $\alpha$ on both the convergence rate and consistency.
@@
 Analogous to Corollary 2, it can be established that if the sparsity level and dimensionality satisfy $k_{t}\sqrt{k_{t}\log p}=o(1)$, the estimation errors for both in-sample and out-of-sample Sharpe ratios during reallocation periods converge to zero as $n\to\infty$.
+Since the conclusion is identical to that of Corollary 2, we omit the formal statement here.
 3 Simulation
@@
 $R_{i}=b_{i1}f_{i1}+b_{i2}f_{i2}+b_{i3}f_{i3}+\varepsilon_{i},\quad i=1,\cdots,p.$
-First, the factor loadings $\bm{b}_{i}$ are drawn from a trivariate normal distribution $\mathsf{N}\left(\bm{\mu}_{b},\mathbf{cov}_{b}\right)$ with parameters specified in Table 1 below.
+First, the factor loadings $\mathbf{b}_{i}$ are drawn from a trivariate normal distribution ${N}\left({\bm{\mu}}_{b},\mathbf{cov}_{b}\right)$ with parameters specified in Table 1 below.
 Once generated, these factor loadings remain fixed throughout the simulations.
-The returns of the three factors over $n$ periods are also sampled from a trivariate normal distribution $\mathsf{N}\left(\bm{\mu}_{f},\mathbf{cov}_{f}\right)$, with the relevant parameters provided in Table 1.
-Throughout this simulation study, we assume that $\mathbb{E}(\mathbf{\bm{\epsilon}}\mid\mathbf{f})=\mathbf{0}$ and $\mathsf{Cov}(\bm{\epsilon}\mid\mathbf{f})=\operatorname{diag}\left(\sigma_{1}^{2},\cdots,\sigma_{p}^{2}\right)$.
+The returns of the three factors over $n$ periods are also sampled from a trivariate normal distribution ${N}\left({\bm{\mu}}_{f},\mathbf{cov}_{f}\right)$, with the relevant parameters provided in Table 1.
+Throughout this simulation study, we assume that $\mathbf{E}(\bm{\varepsilon}\mid\mathbf{f})=\mathbf{0}$ and $\mathbf{Cov}(\bm{\varepsilon}\mid\mathbf{f})$ is a diagonal matrix with entries $\sigma_{1}^{2},\dots,\sigma_{p}^{2}$ on the diagonal.
 The standard deviations of the idiosyncratic errors $\varepsilon_{i}$ are generated from a gamma distribution with a shape parameter of 3.3586 and a scale parameter of 0.1876.
@@
 0.020714 -0.20419 -0.0022526 0.19303
-To examine the performance of the proposed method, we include the following three models for comparison in the simulation study.
-  • 1.
-    The conventional Markowitz mean-variance model
-    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1},\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.1)
-    where the $w_{1,i}$ is denoted as the $i$th element of $\bm{w}_{1}$.
-  • 2.
-    The regularized mean-variance model with $\ell_{1}$ penalty.
-    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{t}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\lambda\|\bm{w}_{1}\|_{1},\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.2)
-  • 3.
-    The mean-variance model with the transaction cost.
-    $\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\mathbf{C}_{1}(\bm{w}_{1}),\quad\text{ s.t. }\sum_{i=1}^{p}w_{1,i}=1.$ (3.3)
-The models described above are for comparison during the first portfolio construction stage.
-By replacing $\bm{w}$ with $\bm{\delta}$ and substituting the subscript $1$ with $t$, the corresponding three models for the reallocation stage can be obtained.
 We set $p=2000$, $m=5$ and $n=200$, dividing the data into five stages to calculate the out-of-sample annual Return, Cost, Turnover, Leverage and Sharpe ratio.
+And S0 is the portfolio construction stage, S1-S4 are rebalancing stages, respectively.
 The optimal tuning parameter $\lambda^{opt}$ is selected based on the highest in-sample Sharpe ratio.
-The main method is the proposed Cost-Aware Portfolio Estimator (CAPE).
-For comparison, we utilize three different baseline models: the Mean-Variance (MV) model (3.1), the Penalized Mean-Variance model (PMV) (3.2), and the Cost-Aware Mean-Variance (CMV) model (3.3) to compute portfolio weights at each stage.
-We employ the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE) used in Ledoit and Wolf (2017) to estimate the covariance matrix $\bm{\Sigma}$.
-Specifically, we utilize the functions nonlin_shrinkLW() and cov.shrink() from the R packages HDShOP and corpcor to calculate NLSE and LSE, respectively.
-The results from the NLSE will be presented in the main text, while the results from the LSE are included in the Supplementary Materials.
+The main method is the proposed Cost-Aware Portfolio Estimator under SCAD Penalty (CAPE-S).
+For comparison, we utilize three different baseline models montioned in Section 2.1: the classical Mean-Variance (MV) model (2.1), the Penalized Mean-Variance (PMV) model (2.2), and the Cost-Aware Mean-Variance (CMV) model (2.3) to compute portfolio weights at each stage.
+We estimate the covariance matrix $\bm{\Sigma}_{t}$ using the linear shrinkage estimator (LSE; Ledoit and Wolf (2004)) and the nonlinear shrinkage estimator (NLSE; Ledoit and Wolf (2020)), while the mean vector $\bm{\mu}_{t}$ is estimated by the sample mean.
+The results based on the LSE, which satisfy Assumptions (A2) and (A3) outlined earlier, are presented in the main text.
+In contrast, the results based on the NLSE, which satisfy only Assumption (A3) but not (A2), are reported in the Supplementary Materials.
+Despite not meeting (A2), the NLSE performs well in both simulation studies and real data applications.
 Table 2: Simulation results with quadratic transaction costs, means and standard errors are calculated based on 1000 replicates, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
 Stage Method Return Cost Turnover Leverage SR
-S2 MV 8.158(0.273) 3.370(0.004) 16.494(0.010) 7.747(0.005) 1.004(0.034)
-PMV 7.911(0.139) 0.740(0.002) 4.363(0.006) 1.681(0.003) 2.020(0.036)
-CMV 8.330(0.244) 2.693(0.003) 14.740(0.008) 6.870(0.004) 1.162(0.034)
-CAPE 7.881(0.128) 0.616(0.002) 3.855(0.005) 1.428(0.003) 2.196(0.036)
-S3 MV 4.138(0.273) 6.482(0.024) 22.684(0.040) 7.754(0.005) 0.436(0.029)
-PMV 8.107(0.162) 1.056(0.004) 5.275(0.009) 2.799(0.003) 1.747(0.035)
-CMV 9.634(0.246) 4.683(0.015) 19.265(0.029) 6.667(0.004) 1.201(0.031)
-CAPE 8.189(0.152) 0.839(0.003) 4.516(0.008) 2.458(0.003) 1.908(0.035)
-S4 MV 4.285(0.275) 6.541(0.024) 22.791(0.040) 7.736(0.005) 0.451(0.029)
-PMV 8.472(0.173) 1.193(0.004) 5.709(0.011) 3.395(0.003) 1.728(0.036)
-CMV 11.748(0.247) 4.564(0.014) 19.036(0.027) 6.527(0.004) 1.476(0.031)
-CAPE 8.554(0.162) 0.950(0.004) 4.882(0.009) 3.034(0.003) 1.874(0.036)
-S5 MV 4.397(0.271) 6.517(0.024) 22.747(0.041) 7.742(0.005) 0.466(0.029)
-PMV 9.036(0.173) 1.244(0.005) 5.902(0.012) 3.733(0.004) 1.804(0.035)
-CMV 12.730(0.245) 4.477(0.013) 18.860(0.026) 6.476(0.004) 1.609(0.031)
-CAPE 9.137(0.163) 0.994(0.004) 5.057(0.010) 3.369(0.004) 1.954(0.035)
-  • Notes: MV, PMV, CMV and CAPE are mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with nonlinear shrinkage estimated covariance matrices), respectively.
-S1 is the portfolio construction stage, S2-S5 are rebalancing stages, respectively.
-And the values in columns of Return, Cost, Turnover, and Leverage are in percentages.
-SR is the Sharpe ratio post transaction costs.
+S1 MV 7.992(0.241) 4.120(0.005) 18.160(0.010) 8.580(0.005) 1.044(0.032)
+PMV 7.766(0.107) 0.624(0.001) 4.082(0.006) 1.541(0.003) 2.599(0.036)
+CMV 8.285(0.212) 2.993(0.003) 15.693(0.009) 7.347(0.004) 1.279(0.033)
+CAPE-S 7.723(0.097) 0.500(0.001) 3.562(0.005) 1.281(0.002) 2.851(0.036)
+S2 MV 3.581(0.242) 7.736(0.024) 18.161(0.011) 8.580(0.005) 0.374(0.025)
+PMV 8.269(0.125) 0.831(0.002) 4.845(0.008) 2.662(0.003) 2.290(0.035)
+CMV 9.635(0.215) 4.974(0.013) 20.216(0.026) 7.186(0.004) 1.294(0.029)
+CAPE-S 8.247(0.115) 0.631(0.002) 4.088(0.007) 2.291(0.003) 2.506(0.035)
+S3 MV 3.670(0.250) 7.792(0.024) 18.131(0.010) 8.566(0.005) 0.385(0.026)
+PMV 8.785(0.139) 0.934(0.003) 5.234(0.009) 3.305(0.003) 2.269(0.036)
+CMV 11.613(0.220) 4.870(0.012) 20.012(0.025) 7.050(0.004) 1.572(0.030)
+CAPE-S 8.758(0.128) 0.708(0.002) 4.406(0.007) 2.894(0.003) 2.469(0.036)
+S4 MV 3.764(0.245) 7.787(0.025) 18.148(0.010) 8.574(0.005) 0.396(0.026)
+PMV 9.312(0.137) 0.981(0.003) 5.419(0.010) 3.692(0.004) 2.332(0.035)
+CMV 12.530(0.215) 4.797(0.012) 19.855(0.024) 6.997(0.004) 1.705(0.030)
+CAPE-S 9.294(0.128) 0.745(0.002) 4.566(0.008) 3.269(0.003) 2.530(0.035)
+  • Notes: MV, PMV, CMV and CAPE-S are mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with linear shrinkage estimated covariance matrices), respectively.
+S0 is the portfolio construction stage, S1-S4 are rebalancing stages, respectively.
+And the values in columns of Return and Cost represent percentages.
 Table 3: Simulation results with proportional transaction costs, means and standard errors are calculated based on 1000 replicates, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
 Stage Method Return Cost Turnover Leverage SR
-S2 MV 9.879(0.273) 1.649(0.001) 16.494(0.010) 7.747(0.005) 1.284(0.036)
-PMV 8.214(0.139) 0.436(0.001) 4.363(0.006) 1.681(0.003) 2.117(0.036)
-CMV 9.801(0.268) 1.578(0.001) 15.783(0.010) 7.391(0.005) 1.304(0.036)
-CAPE 8.132(0.131) 0.383(0.001) 3.828(0.005) 1.414(0.003) 2.224(0.036)
-S3 MV 8.352(0.271) 2.268(0.004) 22.684(0.040) 7.754(0.005) 1.063(0.035)
-PMV 8.635(0.162) 0.528(0.001) 5.275(0.009) 2.799(0.003) 1.889(0.036)
-CMV 8.448(0.267) 2.171(0.004) 21.713(0.038) 7.610(0.005) 1.092(0.034)
-CAPE 8.572(0.155) 0.455(0.001) 4.551(0.008) 2.449(0.003) 1.970(0.036)
-S4 MV 8.547(0.275) 2.279(0.004) 22.791(0.040) 7.736(0.005) 1.092(0.035)
-PMV 9.094(0.173) 0.571(0.001) 5.709(0.011) 3.395(0.003) 1.890(0.036)
-CMV 8.660(0.270) 2.198(0.004) 21.979(0.039) 7.612(0.005) 1.121(0.035)
-CAPE 8.994(0.166) 0.494(0.001) 4.945(0.010) 3.037(0.003) 1.947(0.036)
-S5 MV 8.639(0.271) 2.275(0.004) 22.747(0.041) 7.742(0.005) 1.105(0.035)
-PMV 9.690(0.173) 0.590(0.001) 5.902(0.012) 3.733(0.004) 1.971(0.035)
-CMV 8.708(0.270) 2.197(0.004) 21.974(0.040) 7.627(0.005) 1.128(0.035)
-CAPE 9.610(0.167) 0.514(0.001) 5.135(0.011) 3.386(0.004) 2.030(0.036)
+S1 MV 10.296(0.241) 1.816(0.001) 18.160(0.010) 8.580(0.005) 1.489(0.035)
+PMV 7.983(0.107) 0.408(0.001) 4.082(0.006) 1.541(0.003) 2.698(0.036)
+CMV 10.169(0.235) 1.729(0.001) 17.293(0.010) 8.147(0.005) 1.513(0.035)
+CAPE-S 7.879(0.099) 0.353(0.001) 3.527(0.005) 1.264(0.003) 2.878(0.036)
+S2 MV 8.833(0.241) 2.484(0.004) 24.840(0.039) 8.580(0.005) 1.245(0.034)
+PMV 8.615(0.125) 0.484(0.001) 4.845(0.008) 2.662(0.003) 2.420(0.035)
+CMV 8.965(0.237) 2.368(0.004) 23.684(0.037) 8.408(0.005) 1.287(0.034)
+CAPE-S 8.471(0.117) 0.410(0.001) 4.097(0.007) 2.277(0.003) 2.552(0.036)
+S3 MV 8.968(0.248) 2.494(0.004) 24.937(0.038) 8.566(0.005) 1.267(0.035)
+PMV 9.196(0.139) 0.523(0.001) 5.233(0.009) 3.305(0.003) 2.414(0.037)
+CMV 9.078(0.245) 2.397(0.004) 23.967(0.038) 8.415(0.005) 1.304(0.035)
+CAPE-S 9.023(0.131) 0.444(0.001) 4.438(0.008) 2.892(0.003) 2.523(0.037)
+S4 MV 9.060(0.244) 2.492(0.004) 24.918(0.040) 8.574(0.005) 1.278(0.035)
+PMV 9.751(0.137) 0.542(0.001) 5.419(0.010) 3.692(0.004) 2.483(0.035)
+CMV 9.142(0.242) 2.399(0.004) 23.986(0.040) 8.432(0.005) 1.310(0.035)
+CAPE-S 9.586(0.130) 0.461(0.001) 4.613(0.008) 3.282(0.003) 2.586(0.036)
   • Note: See notes in Table 2.
-To emulate real-world investment scenarios, we report the performance of portfolios across four stages (S2-S5).
+To emulate real-world investment scenarios, we report the performance of portfolios across four stages (S1-S4).
 The first stage involves forming the initial portfolio, akin to establishing a new fund at its inception.
@@
 The costs in the following tables are therefore interpreted as the percentage of the normalized initial wealth that needs to cover the costs of transactions.
-Additionally, we quantify portfolio turnover and leverage at each stage $k=2,3,4,5$.
+Additionally, we quantify portfolio turnover and leverage at each stage $k=1,2,3,4$.
 The definitions of these metrics are as follows:
-$\text{Turnover}_{k}=\sum\limits_{i=1}^{p}|w_{k,i}-w^{+}_{k-1,i}|,w_{1^{+},i}=0,$
+$\text{Turnover}_{k}=\sum\limits_{i=1}^{p}|w_{k,i}-w^{+}_{k-1,i}|\quad\text{and}\quad w^{+}_{0,i}=0,$
 $\text{Leverage}_{k}=\sum\limits_{i=1}^{p}|\min\{w_{k,i},0\}|$
-The simulation results are summarized in Tables 2 and 3.
-The classic Markowitz mean-variance portfolio exhibits the highest turnover, resulting in the largest transaction costs and the lowest out-of-sample Sharpe ratio.
-In contrast, for both the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE), our proposed method, CAPE, demonstrates the highest Sharpe ratio and the lowest transaction costs.
-This finding underscores the importance of incorporating costs into the optimization problem.
+We evaluate the performance of CAPE-S and three benchmark models under both quadratic and proportional transaction cost settings.
+The results, reported in Tables 2 and 3, demonstrate the consistent superiority of our proposed method in balancing return, risk, and cost-efficiency across all rebalancing stages (S1–S4).
+Under quadratic transaction costs (Table 2), CAPE-S yields the highest Sharpe ratio at every stage, significantly outperforming the traditional Mean-Variance (MV), Penalized Mean-Variance (PMV), and Cost-aware Mean-Variance (CMV) models.
+For example, in stage S1, CAPE-S achieves a Sharpe ratio of 2.851, compared to 2.599 (PMV), 1.279 (CMV), and 1.044 (MV).
+Notably, CAPE-S accomplishes this with the lowest transaction cost (0.500%) and lowest turnover (3.562%), indicating efficient reallocation with minimal trading.
+This pattern persists through stages S2 to S4, with CAPE-S consistently attaining higher return-to-risk ratios, while incurring substantially less cost and leverage than the MV and CMV models.
+Similar conclusions hold under proportional transaction costs (Table 3).
+CAPE-S again delivers the best Sharpe ratio in each stage, with performance gains most prominent in S1 and S4, where the Sharpe ratios reach 2.878 and 2.586, respectively.
+Although PMV also exhibits cost-awareness, it sacrifices return and remains less effective in capturing the full benefit of transaction cost control.
+Compared to CMV and MV, which often yield higher raw returns, CAPE-S achieves better overall risk-adjusted performance by avoiding excessive leverage and unnecessary trading.
+For instance, in S2, CAPE-S reduces the turnover to 4.097% and the cost to 0.410%, about 16.5% of the MV model’s transaction cost, while increasing the Sharpe ratio by approximately 105%.
+These findings highlight the advantage of our model in adapting to real-world trading frictions, demonstrating its ability to attain high returns while tightly controlling costs and maintaining portfolio sparsity.
+CAPE-S successfully balances the trade-off between return and cost, validating its practical value for large-scale portfolio management.
 4 Real Data Application
+In this section, we assess the empirical performance of the proposed CAPE-S method using two major U.S.
+equity universes: the S&P 500 and the Russell 2000 indices.
+For each index, we collect daily return data over the investment horizon from 02-Jan-2017 to 23-Dec-2020, and implement portfolio rebalancing every 251 trading days.
+This results in three decision points and four stages: an initial formation period (S0), followed by three rebalancing stages (S1–S3).
+The out-of-sample evaluation period spans from 01-Jan-2018 to 23-Dec-2020.
+A key feature of this empirical exercise is to use real transaction cost data instead of the pre-determined constant costs coefficients.
+First, to ensure consistency, we cleaned the data by removing stocks with incomplete return histories or missing auxiliary data required for real transaction cost estimation (e.g., bid-ask spreads, trading volumes).
+As a result, we retain 457 stocks from the S&P 500 and 935 stocks from the Russell 2000 for analysis.
+The sample of stocks remain broadly representative of their respective index compositions.
+At each decision point, the covariance matrix $\bm{\Sigma}_{t}$ is estimated using two methods: the linear shrinkage estimator (LSE) and the nonlinear shrinkage estimator (NLSE), both based on the most recent 251 trading days.
+Similar to the setting described in Section 3, we report the results based on the LSE in the main text, while the results obtained from the NLSE are provided in Section 3 of the Supplementary Materials.
+Our evaluation includes the CAPE-S method, three benchmark approaches—mean-variance (MV), penalized mean-variance (PMV), and cost-aware mean-variance (CMV)—and the equally-weighted “1/N” portfolio (DeMiguel et al., 2009).
+We compare their performance using several metrics: annual return, transaction cost, turnover, leverage, stage-specific Sharpe ratios (S1–S3), and the overall Sharpe ratio across the entire out-of-sample period.
+In contrast to the simulation studies, where transaction cost coefficients were fixed across assets (i.e., 0.15 for quadratic costs and 0.001 for proportional costs), such fixed cost specification is likely unrealistic in real-world applications.
+In empirical settings, transaction costs often vary substantially across stocks due to liquidity, trading frequency, and other microstructure factors.
+Therefore, we adopt asset-specific transaction cost coefficients for both proportional and quadratic forms, as discussed in the methodology sections.
+For proportional transaction cost coefficients, we follow the empirical estimates provided in Renò et al.
+(2025), who use the bid-ask spread estimator proposed by Hasbrouck (2009) to proxy the proportional transaction costs at the individual stock level.
+These estimates are averaged over the formation window and are available for a wide range of U.S.
+equities from 1927 to the present, including both S&P 500 and Russell 2000 constituents.
+For quadratic transaction cost coefficients, we define the asset-specific cost coefficient as twice the square of the corresponding proportional cost estimate.
+This specification ensures consistency in scaling and penalizes large rebalancing moves more heavily for less liquid stocks.
+Moreover, it reflects the increasing marginal cost nature of market impact, thereby capturing a more realistic trading environment.
+By incorporating heterogeneity in cost parameters, our empirical analysis more accurately reflects the practical implementation of portfolio strategies under realistic trading frictions.
 4.1 S&P 500 Index
-We collected daily S&P 500 Index composition stocks data covering the holding period from 02-Jan-2017 to 23-Dec-2020.
-Our portfolio is rebalanced every 251 trading days, resulting in three decision days and four stages S1-S4 (construction, and then three rebalances).
-The out-of-sample period extends from 01-Jan-2018 to 23-Dec-2020.
-On each decision day, the covariance matrix $\bm{\Sigma}$ is estimated using two different methods: the nonlinear shrinkage estimator (NLSE) and the linear shrinkage estimator (LSE), both based on the most recent 251 daily returns.
-In the main text we show the results using NLSE.
-The similar results obtained using the LSE are presented in Section S3 of the supplementary materials.
-In the study we consider those stocks with complete records over the holding period.
-Specifically, on the first decision day, a stock has a complete set of 251 days of historical returns, and by the last decision day, it also possesses a full set of returns for the subsequent 251 days.
-As a result, our stock pool comprises 462 stocks.
-In addition to our proposed CAPE method and the three comparison methods (MV, PMV and CMV) shown in the simulation studies, we also include the benchmark equally-weighted portfolio (DeMiguel et al., 2009), where the weights are given as $1/p$.
-Following the convention of the literature we just label it as ‘1/N’ portfolio in the empirical tables.
-We assess several performance metrics, including out-of-sample annual return, transaction costs, turnover, leverage, out-of-sample risk, the annualized Sharpe ratio (we report it separately for stages S2-S4) for holding the portfolio at its current rebalanced composition until the next rebalancing, and the overall Sharpe ratio during the whole S2-S4 period.
-The performances of the various methods are reported in Tables 4 and 5, which present results for proportional and quadratic transaction costs, respectively.
-Our proposed CAPE method demonstrates the best overall Sharpe ratio performance during the S2 to S4 phases.
-In terms of annual returns, the CAPE method has shown the ability to control losses during the period (S2) of overall declines in the S&P 500 and maintaining substantial gains at the lower level of risk in the period of overall profitability (S3-S4).
-Excluding the equally-weighted (‘1/N’) portfolio, the CAPE method exhibits one of the lowest turnover among actively managed portfolios.
-More importantly, CAPE strikes a balance between the required return and risk, considering transaction costs.
-The CAPE is usually the best performer in annual Sharpe ratio, and has the best Sharpe ratio in overall out-of-sample period S2-S4.
-Furthermore, the CAPE method is more robust than the other comparison methods concerning different transaction cost functions.
-Table 4: S&P 500 quadratic transaction costs, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
+We first present the empirical results on the S&P 500 dataset.
+The performance of all methods under both quadratic and proportional transaction cost models is summarized in Tables 4 and 5, respectively.
+Table 4: S&P 500 with quadratic transaction cost.
 Method Stage Return Cost Turnover Leverage SR Overall SR
-1/N S2 -3.835 0.033 1.000 0.000 -0.244 0.863
-S3 29.247 0.002 0.181 0.000 2.336
-S4 47.154 0.002 0.153 0.000 1.067
-MV S2 -2.797 1.034 4.382 1.691 -0.178 0.883
-S3 29.977 2.474 6.754 2.148 2.404
-S4 44.761 1.878 5.878 1.755 1.052
-PMV S2 -6.943 0.826 3.550 1.275 -0.466 0.704
-S3 24.586 0.796 1.859 1.530 2.387
-S4 37.516 0.291 0.780 1.460 0.905
-CMV S2 -2.189 0.808 3.874 1.437 -0.145 0.755
-S3 42.578 3.109 7.611 2.799 2.614
-S4 46.451 2.603 6.920 2.385 0.742
-CAPE S2 -2.016 0.818 3.444 1.222 -0.133 0.970
-S3 30.755 1.183 3.226 1.626 2.822
-S4 57.831 0.588 1.842 1.433 1.204
-  • Notes: ‘1/N’, MV, PMV, CMV and CAPE are the equally-weighted, mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with nonlinear shrinkage estimated covariance matrices), respectively.
-S1 is the portfolio construction stage, S2-S4 are rebalancing stages, respectively.
+1/N S1 -3.666 0.017 1.000 0.000 -0.234 0.869
+S2 29.236 0.001 0.181 0.000 2.341
+S3 47.369 0.002 0.154 0.000 1.073
+MV S1 -7.313 2.545 9.794 4.397 -0.329 0.772
+S2 36.261 13.795 19.462 7.166 1.254
+S3 141.735 12.551 18.118 6.441 1.158
+PMV S1 -10.570 1.346 6.082 2.541 -0.570 0.665
+S2 23.493 1.569 3.081 2.867 2.150
+S3 78.677 0.532 1.561 2.796 1.027
+CMV S1 -5.150 1.488 7.536 3.268 -0.266 0.891
+S2 48.870 8.571 15.235 6.338 2.081
+S3 96.660 6.899 13.651 5.504 1.127
+CAPE-S S1 -3.687 1.631 6.951 2.976 -0.188 0.915
+S2 30.491 2.349 4.766 3.265 2.490
+S3 77.174 0.778 2.283 2.873 1.257
+  • Notes: ‘1/N’, MV, PMV, CMV and CAPE are the equally-weighted, mean variance, penalized mean variance, cost-aware mean variance, and the proposed cost-aware portfolios (all with linear shrinkage estimated covariance matrices), respectively.
+S1 is the portfolio construction stage, S2–S3 are rebalancing stages, respectively.
 And the values in columns of Return and Cost represent percentages.
-The “overall SR” in the last column is the post-transaction cost Sharpe ratio of the portfolio for holding it from S2-S4 to show the overall performance in our sample period.
-More details of other performance measure can be found in the notes of Table 2.
-Table 5: S&P 500 proportional transaction costs, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
+Table 5: S&P 500 with proportional transaction cost.
 Method Stage Return Cost Turnover Leverage SR Overall SR
-1/N S2 -3.903 0.100 1.000 0.000 -0.249 0.862
-S3 29.230 0.018 0.181 0.000 2.335
-S4 47.141 0.015 0.153 0.000 1.067
-MV S2 -2.201 0.438 4.382 1.691 -0.140 0.930
-S3 31.775 0.675 6.754 2.148 2.603
-S4 46.051 0.588 5.878 1.755 1.083
-PMV S2 -6.472 0.355 3.550 1.275 -0.434 0.852
-S3 31.187 0.622 6.224 2.140 2.598
-S4 58.348 0.535 5.354 1.719 1.103
-CMV S2 -7.849 0.429 4.295 1.647 -0.518 0.844
-S3 29.955 0.707 7.067 2.299 2.514
-S4 63.214 0.634 6.339 1.862 1.148
-CAPE S2 -2.270 0.347 3.467 1.234 -0.149 1.000
-S3 31.757 0.641 6.410 2.288 2.519
-S4 52.784 0.578 5.783 1.802 1.223
+1/N S1 -3.797 0.147 1.000 0.000 -0.242 0.866
+S2 29.185 0.053 0.181 0.000 2.336
+S3 47.323 0.047 0.154 0.000 1.072
+MV S1 -6.092 1.324 9.794 4.397 -0.274 0.859
+S2 44.854 5.203 19.462 7.166 1.730
+S3 149.530 4.756 18.118 6.441 1.228
+PMV S1 -10.058 0.833 6.082 2.541 -0.543 0.824
+S2 37.566 3.902 14.298 6.180 1.721
+S3 134.396 3.730 13.927 5.555 1.225
+CMV S1 -10.864 1.133 8.474 3.737 -0.523 0.798
+S2 38.339 4.626 17.411 6.993 1.612
+S3 154.734 4.345 16.763 6.321 1.209
+CAPE-S S1 -2.694 0.949 7.147 3.074 -0.133 0.909
+S2 37.775 4.128 15.688 6.606 1.666
+S3 108.519 3.912 15.051 6.007 1.263
   • Note: See notes in Table 4.
+Our proposed CAPE-S method achieves the highest overall out-of-sample Sharpe ratio across the three stages (S1 to S3) under both quadratic and proportional transaction cost settings.
+Examining performance at the stage level provides further insights:
+  • Stage S1 (2018) coincided with a period of heightened market volatility and an overall downturn, particularly during Q4 2018 when the S&P 500 experienced a sharp correction.
+While all active strategies posted negative returns, CAPE-S recorded the smallest loss among them under the proportional cost setting (-2.694%) and one of the smallest losses under the quadratic cost setting (-3.687%), demonstrating resilience during stressed market conditions.
+Turnover (7.147% proportional, 6.951% quadratic) and leverage (3.074 proportional, 2.976 quadratic) remained moderate relative to other active strategies.
+  • Stage S2 (2019) represented a strong recovery year, with broad-based gains across the S&P 500 driven by accommodative monetary policy and easing trade tensions.
+Under both cost structures, CAPE-S delivered competitive returns (37.775% proportional, 30.491% quadratic) and achieved one of the highest Sharpe ratios among actively managed portfolios (1.666 proportional, 2.490 quadratic), while maintaining transaction costs (4.128% proportional, 2.349% quadratic) and turnover (15.688% proportional, 4.766% quadratic) at moderate levels compared to MV, PMV, and CMV.
+  • Stage S3 (2020) covered a highly turbulent market cycle, including the COVID-19-induced crash in Q1 2020 followed by a rapid rebound.
+CAPE-S generated solid returns (108.519% proportional, 77.174% quadratic) and achieved the highest Sharpe ratio under the proportional cost setting (1.263), while keeping leverage (6.007 proportional, 2.873 quadratic) relatively contained compared to MV and CMV.
+This highlights its adaptability to sharp market reversals.
+Across all stages, excluding the passive 1/N portfolio, CAPE-S consistently maintains one of the lowest turnover levels among actively managed strategies, reflecting operational efficiency.
+Its leading overall Sharpe ratio under both cost specifications (0.915 quadratic, 0.909 proportional) highlights its capability to balance returns, risk, and transaction costs, demonstrating robustness to varying cost structures.
 4.2 Russel 2000 Index
-In this subsection we evaluate the performance of CAPE in the relatively small-cap stocks pool.
-We collected daily Russell 2000 Index composition stocks data for the holding period from 02-Jan-2017 to 23-Dec-2020.
-Our portfolio is rebalanced every 251 trading days, resulting in three decision days and four stages.
-The out-of-sample period extends from 01-Jan-2018 to 23-Dec-2020.
-We employ the same criterion of complete observations described in Section 4.1, resulting in a pool of 991 stocks.
-This scenario imposes great challenges for portfolio management in practice, however, it showcases the advantage and applicability of our proposed method.
-The performance of the various methods is reported in Tables 6 and 7, which correspond to proportional and quadratic transaction costs, respectively.
-The performances of MV and CMV methods are notably poor, as demonstrated in Tables 6 and 7.
-The PMV method exhibits a slightly better overall Sharpe ratio during the S2 to S4 phases compared to the MV and CMV methods.
-In contrast, our CAPE method significantly outperforms the other approaches, achieving the highest overall Sharpe ratio.
-The turnover ratio of CAPE is notably the smallest among all actively managed portfolios.
-Consistent with our findings in the S&P 500 large-cap stocks, the CAPE method demonstrates robustness to variations in transaction cost functions.
-Table 6: Russell 2000 quadratic transaction costs, the unit rebalancing cost of the portfolio $\beta=0.15$ and $\gamma=1/3$.
+Next, we report the results for the Russell 2000 dataset, following the same setup as in Section 4.1. The corresponding performance metrics are presented in Tables 6 and 7.
+Table 6: Russell 2000 with quadratic transaction cost.
 Method Stage Return Cost Turnover Leverage SR Overall SR
-1/N S2 -8.008 0.015 1.000 0.000 -0.504 0.876
-S3 44.196 0.003 0.285 0.000 1.810
-S4 140.395 0.015 0.287 0.000 1.247
-MV S2 -1.192 0.439 4.176 1.588 -0.080 1.010
-S3 133.689 0.706 5.013 1.598 1.113
-S4 106.914 1.248 5.681 1.394 1.674
-PMV S2 3.114 0.193 1.802 0.401 0.286 1.111
-S3 114.843 0.183 1.607 0.713 1.224
-S4 102.532 0.332 1.728 0.706 1.583
-CMV S2 -0.946 0.413 4.049 1.524 -0.065 1.060
-S3 189.224 0.778 5.402 1.716 1.144
-S4 188.827 1.518 6.000 1.565 1.562
-CAPE S2 8.044 0.154 1.194 0.097 0.865 1.122
-S3 116.971 0.089 0.827 0.344 1.278
-S4 88.369 0.224 0.833 0.351 1.474
+1/N S1 -7.839 0.012 1.000 0.000 -0.488 0.816
+S2 25.682 0.003 0.281 0.000 1.656
+S3 124.503 0.017 0.281 0.000 1.269
+MV S1 -55.130 6.625 19.522 9.261 -1.323 0.518
+S2 -25.690 23.272 34.829 9.993 -0.406
+S3 261.635 23.275 34.694 11.870 1.404
+PMV S1 -10.425 0.566 2.762 0.881 -0.686 1.243
+S2 21.163 1.216 4.343 2.144 0.987
+S3 122.167 1.880 5.011 2.788 2.197
+CMV S1 -46.062 4.543 16.201 7.600 -1.290 0.784
+S2 12.367 12.775 25.798 8.301 0.237
+S3 366.413 12.120 25.018 9.537 1.551
+CAPE S1 3.273 0.319 1.307 0.153 0.297 1.170
+S2 32.691 0.280 1.519 0.695 2.745
+S3 177.988 0.356 1.600 0.973 1.710
   • Note: See notes in Table 4.
-Table 7: Russell 2000 proportional transaction costs, the unit rebalancing cost of the portfolio $\alpha=0.001$ and $\gamma=1/3$.
+Table 7: Russell 2000 with proportional transaction cost.
 Method Stage Return Cost Turnover Leverage SR Overall SR
-1/N S2 -8.093 0.100 1.000 0.000 -0.510 0.876
-S3 44.171 0.029 0.285 0.000 1.809
-S4 140.381 0.029 0.287 0.000 1.247
-MV S2 -1.170 0.418 4.176 1.588 -0.079 1.014
-S3 133.893 0.501 5.013 1.598 1.115
-S4 107.594 0.568 5.681 1.394 1.685
-PMV S2 3.127 0.180 1.802 0.401 0.287 1.112
-S3 114.865 0.161 1.607 0.713 1.224
-S4 102.691 0.173 1.728 0.706 1.585
-CMV S2 -3.548 0.439 4.391 1.695 -0.249 1.028
-S3 141.701 0.527 5.265 1.735 1.081
-S4 126.576 0.745 7.454 2.066 1.841
-CAPE S2 8.104 0.119 1.190 0.095 0.872 1.128
-S3 116.553 0.082 0.824 0.341 1.280
-S4 89.496 0.083 0.833 0.349 1.481
+1/N S1 -8.157 0.330 1.000 0.000 -0.508 0.813
+S2 25.569 0.116 0.281 0.000 1.649
+S3 124.399 0.121 0.281 0.000 1.268
+MV S1 -54.555 6.050 19.522 9.261 -1.311 0.599
+S2 -13.976 11.559 34.829 9.993 -0.240
+S3 274.625 10.285 34.694 11.870 1.484
+PMV S1 -10.945 1.086 2.762 0.881 -0.723 1.234
+S2 20.627 1.752 4.343 2.144 0.961
+S3 122.238 1.809 5.011 2.788 2.199
+CMV S1 -57.093 5.355 17.963 8.482 -1.465 0.533
+S2 -47.534 10.907 33.960 10.564 -0.849
+S3 483.051 21.687 75.347 17.964 1.196
+CAPE S1 2.603 0.594 1.282 0.141 0.237 1.158
+S2 32.209 0.678 1.477 0.665 2.732
+S3 177.379 0.635 1.568 0.956 1.699
   • Note: See notes in Table 4.
+Despite the higher volatility and liquidity challenges inherent in small-cap stocks, the CAPE-S method maintains robust performance under both quadratic and proportional transaction cost settings.
+Tables 6 and 7 report detailed results under the two cost regimes.
+  • Stage S1 (2018) was a period of elevated volatility and broad market drawdown.
+The Russell 2000 Index, which is generally more sensitive to macroeconomic uncertainty, suffered larger losses than the S&P 500.
+In this challenging environment, CAPE-S was the only strategy to achieve positive returns, with modest gains accompanied by low turnover (approximately 1.3%) and low leverage (around 0.15), highlighting its defensiveness and cost-efficiency.
+  • Stage S2 (2019) was characterized by a broad market rebound, driven by easing monetary policy and reduced trade tensions.
+During this recovery phase, CAPE-S achieved the highest Sharpe ratio (2.745) among all strategies, while maintaining the lowest transaction cost (0.280%) and turnover (approximately 1.52%), far below the MV strategy’s cost of 23.272% and turnover of 34.829%.
+This demonstrates that CAPE-S effectively balances risk and return without relying on aggressive rebalancing.
+  • Stage S3 (2020) encompassed one of the most turbulent and rapidly rebounding market periods in recent history due to the COVID-19 pandemic.
+The small-cap segment staged a dramatic recovery following the March 2020 crash, benefiting disproportionately from fiscal stimulus and risk-on sentiment.
+CAPE-S captured substantial gains (approximately 177.99%) while preserving a favorable risk profile, with low leverage (0.973) and turnover (1.60%), outperforming all alternatives in terms of Sharpe ratio (1.710).
+Overall, the MV and CMV strategies performed poorly, particularly under proportional cost settings, due to high turnover and leverage.
+The PMV method showed some improvement but still lagged behind CAPE-S.
+Across all three stages, CAPE-S consistently achieved the highest overall Sharpe ratio (1.170 for quadratic costs and 1.158 for proportional costs) while maintaining the lowest turnover among actively managed portfolios.
+These findings confirm CAPE-S’s adaptability and robustness, particularly in the more volatile and cost-sensitive small-cap universe, and underscore its practical relevance for real-world portfolio management.
 5 Conclusion
-In this paper, we propose a novel approach to integrate the transaction cost within the optimization problem when the number of assets is larger than the sample size.
-By using LLA algorithm, we prove the sign consistency and provide the $\ell_{\infty}$ bound of our proposed CAPE method for both proportional and quadratic transaction costs.
-In addition to the sound statistical properties, CAPE demonstrates exceptional performance in practical applications.
-Through extensive simulations and applications to real data we demonstrate the superior performance and broad applicability of our method across diverse tasks.
+In this paper, we propose a novel approach to integrate the transaction costs within the optimization problem when the number of assets is larger than the sample size.
+By using LLA algorithm, we prove the sign consistency and provide the $\ell_{\infty}$ bound of our proposed CAPE-S method for both proportional and quadratic transaction costs.
+In addition to the sound statistical properties, CAPE-S demonstrates exceptional performance in practical applications.
+Through extensive simulations and applications to real data, we demonstrate the superior performance and broad applicability of our method across diverse tasks.